### v1

# A Visual Approach to the Regularity Theory for Fully Nonlinear Elliptic Equations

Abstract.
In this set of notes we revisit the regularity theory for uniformly elliptic equations without divergence structure.
Key words and phrases:
Elliptic regularity theory, Krylov-Safonov, Harnack inequalities.
1991 Mathematics Subject Classification:
(Primary) 35-01, (Secondary) 35B45, 35B65, 35J05, 35J15, 35J60, 35J70, 35K10, 35K55, 31B05
1. Introduction
The analysis of partial differential equations (PDEs) is a well-established discipline that treasures beautiful geometric insights. While these concepts often come to life during seminar talks or discussions with experts in the field, students approaching these topics individually may encounter difficulties in extracting geometric meaning from the typical computations found in textbooks and research articles.
In these notes, we aim to provide a pedagogical perspective that places a strong emphasis on the visual approach. Our journey begins with the fundamental case, that of the Laplacian, and progressively leads to the Krylov-Safonov regularity theory. Along the way, we engage the reader with a series of challenging problems providing a valuable opportunity for hands-on learning. These problems are far from easy, many of them were actually taken from results in contemporary research articles, however we expect that the techniques presented give some clues on their solutions. Towards the end, we offer hints into how to tackle these problems and provide references for those seeking a more comprehensive treatment. Finally, we also offer some discussions on some further developments and the some open problems in the field.
The main idea we hope to illustrate is how uniform ellipticity enforces regularity on the solutions of elliptic equations through measure estimates. The main example of this phenomenon is the mean value property for harmonic functions. In the next section we will see how it implies an interior Hölder estimate on any harmonic function, depending exclusively on its oscillation.
To appreciate the significance of this result, consider an arbitrary uniformly bounded sequence of functions defined over a compact set. Notice that, in contrast to numerical sequences, a bounded sequence of functions does not necessarily possess a convergent sub-sequence, even when each function within the sequence is smooth. However, as soon as we assume that the functions are harmonic, then such limits are always guaranteed in the sense of uniform convergence by the Arzelà–Ascoli theorem and the regularity estimate for harmonic functions. This compactness plays an important role in numerous fundamental results in analysis of elliptic PDEs, such as existence theorems or the convergence of numerical schemes.
The analysis of PDEs heavily relies on establishing bounds for the modulus of continuity of a solution and its derivatives. These estimates serve as the cornerstones of what is known as the regularity theory.
The development of the regularity theory for elliptic equations has a rich history, with numerous contributing authors. These notes delves into the regularity theory for uniformly elliptic equations without divergence structure, a field that originated during 1980s and 1990s together with the theory of viscosity solutions. This period witnessed significant contributions from notable figures such as Krylov, Safonov, Evans, Trudinger, and Caffarelli, among many others. Collectively, their work has given rise to what is now widely recognized as the Krylov-Safonov Theory. For a thorough discussion on the history of the subject we recommend [BB98, Kry97], Chapter 1 in [NTV14], and the notes at the end of Chapter 9 in [GT83].
We do not provide in this note a discussion about the rich variety of problems that motivate the study of elliptic equations. For this, we recommend the recent manuscript [DV23], where a wide array of intriguing real-world problems, ranging from heat conduction, population dynamics, and random walks are explored in depth. The book [Mar07] provides a visual approach discussing applications for broader range of PDEs.
The subjects we present in this article have been covered in various books [GT83, CC95, Kry96, HL11, FRRO22], notes [Moo12], and articles [Wan92, Cab97, Sil06, Sav07, IS16, Moo15].
These notes originated from a lecture offered by the author at the Learning Seminar on Analysis of PDEs for the Centre for Mathematics at the University of Coimbra on July of 2021. They are dedicated to Luis Caffarelli, with profound gratitude and admiration.
Acknowledgment: The author was supported by CONACyT-MEXICO grant A1-S-48577.
2. The Laplacian
In this section, we demonstrate how the most basic form of uniform ellipticity leads to interior Hölder estimates. The main result is stated in Theorem 2.3, which will be derived from the diminish of oscillation strategy in Lemma 2.6.
In the following results, we use the oscillation of a function to measure the variation of the values that it takes in a given set. For $u\colon\Omega\subseteq{\mathbb{R}}^{n}\to{\mathbb{R}}$ and $E\subseteq\Omega$,
$\operatorname*{osc}_{E}u:=\sup_{x,y\in E}|u(y)-u(x)|=\sup_{E}u-\inf_{E}u.$
For a fixed $x_{0}\in\Omega$, the oscillation of $u$ over [^1] We will constantly use the following notation for open balls $B_{r}(x_{0}):=\{x\in{\mathbb{R}}^{n}\,:\,|x-x_{0}|<r\}$. By default $B_{r}:=B_{r}(0)$. $B_{r}(x_{0})\cap\Omega$ can be used to determine the continuity of $u$ around $x_{0}$. Indeed, $u$ is continuous at $x_{0}$ if and only if
$\lim_{r\to 0^{+}}\operatorname*{osc}_{B_{r}(x_{0})\cap\Omega}u=0.$
Moreover, $u$ is uniformly continuous if and only if
$\lim_{r\to 0^{+}}\sup_{x_{0}\in\Omega}\operatorname*{osc}_{B_{r}(x_{0})\cap\Omega}u=0.$
Any increasing function $\omega\colon[0,\infty)\to[0,\infty)$ such that $\omega(0)=0$ and $\omega(r)\geq\operatorname*{osc}_{B_{r}(x_{0})\cap\Omega}u$, is said to be a modulus of continuity for $u$ at $x_{0}$. If we also have $\omega(r)\geq\sup_{x_{0}\in\Omega}\operatorname*{osc}_{B_{r}(x_{0})}u$ then we just said that it is a modulus of continuity for $u$. The idea is that $\omega$ quantifies the continuity of $u$, either punctually of globally.
Figure 1. An $\alpha$-Hölder continuous function is one whose graph can be squeezed at every point $(x_{0},u(x_{0}))$ between the graphs of $u(x_{0})-C|x-x_{0}|^{\alpha}$ and $u(x_{0})+C|x-x_{0}|^{\alpha}$ for some $C>0$. Equivalently, $\operatorname*{osc}_{B_{r}(x_{0})\cap\Omega}u\simeq\sup_{B_{r}(x_{0})\cap\Omega}|u-u(x_{0})|\leq Cr^{\alpha}$.
One of the most important examples of modulus of continuity is the Hölder modulus, it has the form $\omega(r):=Cr^{\alpha}$ for some $C>0$ and $\alpha\in(0,1]$, see Figure 1. One of the most important features of this modulus of continuity is the scaling symmetry that it possesses which leads to the diminish of oscillation characterization
Lemma 2.1 (Characterization of Hölder Continuity by Diminish of Oscillation).
Consider $\rho,\theta\in(0,1)$, and $u\colon B_{1}\to{\mathbb{R}}$ a bounded function. If $u$ satisfies the following diminish of oscillation identity for every radius $r\in(0,1]$
$\operatorname*{osc}_{B_{\rho r}}u\leq(1-\theta)\operatorname*{osc}_{B_{r}}u,$
then $u$ admits the following Hölder modulus of continuity at the origin
$\operatorname*{osc}_{B_{r}}u\leq Cr^{\alpha}\operatorname*{osc}_{B_{1}}u,$
where $\alpha:=\ln(1-\theta)/\ln\rho$, $C:=(1-\theta)^{-1}$, and $r\in(0,1]$.
We will usually have that $\theta$ and $\rho$ are small positive constants. In particular, $1-\theta\geq\rho$ would make $\alpha\in(0,1]$.
Proof.
By an inductive argument we have that for any integer $k\geq 0$
$\operatorname*{osc}_{B_{\rho^{k}}}u\leq(1-\theta)^{k}\operatorname*{osc}_{B_{1}}u.$
From this identity we can see why $\alpha$ was set to be $\ln(1-\theta)/\ln\rho$. It is the exponent for which we have that $(1-\theta)^{k}=(\rho^{k})^{\alpha}$, for any $k$.
Let $r\in(0,1]$ and $k=\lfloor\ln r/\ln\rho\rfloor\geq 0$ such that $\rho^{k+1}<r\leq\rho^{k}$. Then
$\operatorname*{osc}_{B_{r}}u\leq\operatorname*{osc}_{B_{\rho^{k}}}u\leq(1-\theta)^{k}\operatorname*{osc}_{B_{1}}u=C\rho^{\alpha(k+1)}\operatorname*{osc}_{B_{1}}u\leq Cr^{\alpha}\operatorname*{osc}_{B_{1}}u.$
This concludes the proof. ∎
Remark 2.2.
The previous proof only uses that the diminish of oscillation holds for $r=\rho^{k}$, for any integer $k\geq 0$.
Given $\alpha\in(0,1]$, we measure the Hölder semi-norm of $u\colon\Omega\to{\mathbb{R}}$ over $E\subseteq\Omega$ in the following form
$[u]_{C^{0,\alpha}(E)}:=\sup_{x,y\in E}\frac{|u(y)-u(x)|}{|y-x|^{\alpha}}=\sup_{\begin{subarray}{c}x_{0}\in E\\ r>0\end{subarray}}r^{-\alpha}\operatorname*{osc}_{B_{r}(x_{0})\cap E}u.$
For $\beta\in{\mathbb{R}}$, we denote the corresponding interior weighted semi-norm in $\Omega$ as
$[u]_{C^{0,\alpha}(\Omega)}^{(\beta)}:=\sup_{B_{r}(x_{0})\subseteq\Omega}r^{\beta}[u]_{C^{0,\alpha}(B_{r/2}(x_{0}))}.$
Notice that $[u]_{C^{0,\alpha}(\Omega)}^{(\beta)}$ controls how the $\alpha$-Hölder modulus of $u$ at $x_{0}\in\Omega$ behaves in terms of the distance from $x_{0}$ to the boundary of $\Omega$. Indeed, for $d:=\operatorname*{dist}(x_{0},\partial\Omega)$,
$\sup_{r\in(0,d/2)}r^{-\alpha}\operatorname*{osc}_{B_{r}(x_{0})}u\leq d^{-\beta}[u]_{C^{0,\alpha}(\Omega)}^{(\beta)}.$
So, for $\beta>0$ and $[u]_{C^{0,\alpha}(\Omega)}^{(\beta)}<\infty$, we have that $u$ is locally $\alpha$-Hölder continuous in $\Omega$, and we may allow the constant to degenerate at the boundary with an algebraic rate.
In a similar fashion we extend this notion for arbitrary semi-norms on real valued functions defined over $\Omega$. For instance, the $L^{p}$-norm ($p\geq 1$) [^2] The simple $L^{\infty}$-norm given here is valid for continuous functions, which will be the case in these notes. In the case of measurable and possibly discontinuous functions, one must use instead the least bound for $|u|$ up to sets of Lebesgue measure zero $\|u\|_{L^{\infty}(E)}:=\inf\{M>0\,:\,|\{x\in E\,:\,|u(x)|>M\}|=0\}.$ It would be better to use a different notation for continuous functions, perhaps the uniform norm $\|\cdot\|_{C^{0}}$, however $\|\cdot\|_{L^{\infty}}$ is the standard notation in most references.
$\|u\|_{L^{p}(\Omega)}^{(\beta)}:=\sup_{B_{r}(x_{0})\subseteq\Omega}r^{\beta}\|u\|_{L^{p}(B_{r/2}(x_{0}))},$
$\|u\|_{L^{p}(E)}:=\left(\int_{E}|u|^{p}\right)^{1/p},\qquad(p\in[1,\infty)])$
$\|u\|_{L^{\infty}(E)}:=\sup_{E}|u|.$
We finally have the ingredients to state the main regularity theorem of this section.
Theorem 2.3 (Interior Hölder Estimate - Laplacian).
Given $p>n/2$, there exist $\alpha\in(0,1)$ and $C\geq 1$ such that for $\Omega\subseteq{\mathbb{R}}^{n}$ open, and $u\in C^{2}(\Omega)$ we have that
$[u]_{C^{0,\alpha}(\Omega)}^{(\alpha)}\leq C\left(\operatorname*{osc}_{\Omega}u+\|\Delta u\|_{L^{p}(\Omega)}^{(2-n/p)}\right).$
In the subsequent results, we consistently maintain the dimension of the Euclidean space as $n$. All the constants mentioned in the following discussion will always depend on this dimension, unless stated otherwise.
The first observation about Theorem 2.3 is that, due to the definitions of the weighted semi-norms, we only need to establish a Hölder estimate in balls contained in $\Omega$. Actually, by performing a standard change of variable, we can focus just on functions defined over the unit ball $B_{1}$, with the goal of giving an estimate on the Hölder modulus of continuity at the origin.
The strategy to prove such estimate consists in showing that the oscillation of $u$ has a geometric decay in a sequence of concentric balls with geometrically decreasing radii (cf. Lemma 2.1). To be more specific, we will see that the decay in the oscillation of $u$ from a ball of radius $r$ to a concentric ball of radius $r/3$ is determined by the average of $u$ over the ball of radius $r/3$. The mean value property, a consequence of the divergence theorem, plays a crucial role in enabling this approach.
Lemma 2.4 (Mean Value Property).
Given $p>n/2$, there exists $C\geq 1$ such that for $u\in C^{2}(B_{1})$ we have that [^3] Given $a\in{\mathbb{R}}$, we denote its positive and negative parts respectively as $a_{+}=\max(a,0)$ and $a_{-}=\max(-a,0)$, such that $a=a_{+}-a_{-}$.
$\fint_{B_{1}}u\leq u(0)+C\|(\Delta u)_{+}\|_{L^{p}(B_{1})}.$
Figure 2. Given $\Delta u\leq 0$ in $\Omega$, we get that for any ball $B_{r}(x_{0})$ contained in $\Omega$, the average of $u$ over $B_{r}(x_{0})$ is less than the value at the center.
Proof.
Consider for $r\in(0,1)$, the average of $u$ over $\partial B_{r}$. By the change of variable formula
[…]
Then we compute its derivative with respect to $r$
[…]
The second equality follows by the divergence theorem. We now apply Hölder’s inequality to get
$\frac{d}{dr}\fint_{\partial B_{r}}u\leq Cr^{1-n/p}\|(\Delta u)_{+}\|_{L^{p}(B_{1})}.$
Notice that the singularity on the right-hand side is integrable at $r=0$ if and only if $p>n/2$. By integrating the inequality from $r=0$ to $r=\rho\in(0,1)$
$\fint_{\partial B_{\rho}}u\leq u(0)+C\rho^{2-n/p}\|(\Delta u)_{+}\|_{L^{p}(B_{1})}.$
We finally recover the desired formula by integrating from $\rho=0$ to $\rho=1$. ∎
The previous result is also valid in any ball $B_{r}(x_{0})$, namely
$\fint_{B_{r}(x_{0})}u\leq u(x_{0})+Cr^{2-n/p}\|(\Delta u)_{+}\|_{L^{p}(B_{r}(x_{0}))}.$
This can be shown by applying the lemma to $v(x):=u(rx+x_{0})$.
In the homogeneous case, $(\Delta u)_{+}=0$ or equivalently $\Delta u\leq 0$, it gives a particular case of Jensen’s inequality for concave functions as depicted in Figure 2. Functions that satisfy the inequality $\Delta u\leq 0$ are called super-harmonic, and those that satisfy $\Delta u\geq 0$ are called sub-harmonic. While the naming convention might initially appear counterintuitive given the direction of the inequalities, it is actually based on the geometric characteristics of these functions. Specifically, concave functions are super-harmonic and their graphs curve upwards. Conversely, convex functions are sub-harmonic functions and their graphs curve downwards. See Problem 3 for some further analogies.
If we apply Lemma 2.4 to a non-negative function defined in $B_{1}$, and over any possible ball $B_{2/3}(x_{0})$ with $x_{0}\in B_{1/3}$ we deduce the following result. Figure 3 illustrates the idea in the super-harmonic case. The general case can be recovered with the same argument.
Figure 3. For any $x_{0}\in B_{1/3}$, we always have the inclusions $B_{1/3}\subseteq B_{2/3}(x_{0})\subseteq B_{1}$. We used that $u\geq 0$ in the second inequality.
Lemma 2.5 (Weak Harnack Inequality - Laplacian).
Given $p>n/2$, there exists $C\geq 1$ such that for $u\in C^{2}(B_{1})$ non-negative
$\fint_{B_{1/3}}u\leq C\left(\inf_{B_{1/3}}u+\|(\Delta u)_{+}\|_{L^{p}(B_{1})}\right).$
The following lemma is the key estimate which establishes a Hölder modulus of continuity by iterating the given estimate over balls of radii $3^{-k}$. The hypothesis $\|\Delta u\|_{L^{p}(B_{1})}\leq\delta$, for $\delta>0$ sufficiently small, can be considered as a perturbation of the Laplace equation $\Delta u=0$ in $B_{1}$.
Lemma 2.6 (Diminish of Oscillation - Laplacian).
Given $p>n/2$, there exist $\delta,\theta\in(0,1)$, such that for $u\in C^{2}(B_{1})$ with $\|\Delta u\|_{L^{p}(B_{1})}\leq\delta$ we have that
$\operatorname*{osc}_{B_{1}}u\leq 1\qquad\Rightarrow\qquad\operatorname*{osc}_{B_{1/3}}u\leq(1-\theta).$
Proof.
Assume, without loss of generality, that $0\leq u\leq 1$. We now consider two possible alternatives: either $\fint_{B_{1/3}}u$ is $\geq 1/2$ or $\leq 1/2$. In the former case, we get by the weak Harnack inequality that $\inf_{B_{1/3}}u\geq 1/(2C)-\delta$. Otherwise, we apply the same reasoning to $(1-u)$ to conclude that $\sup_{B_{1/3}}u\leq 1-(1/(2C)-\delta)$. In any case, we get the desired diminish of oscillation if we choose $\delta=\theta=1/(4C)$. ∎
Now the idea is to apply the previous lemma over a family of concentric balls in order to obtain a geometric decay of the oscillation as illustrated by Figure 4 (for some $\rho\in(0,1/3]$).
Lemma 2.7 (Hölder Modulus of Continuity - Laplacian).
Given $p>n/2$, there exist $\alpha\in(0,1)$ and $C\geq 1$ such that for $u\in C^{2}(B_{1})$
$\sup_{r\in(0,1)}r^{-\alpha}\operatorname*{osc}_{B_{r}}u\leq C\left(\operatorname*{osc}_{B_{1}}u+\|\Delta u\|_{L^{p}(B_{1})}\right).$
Figure 4. Diminish of oscillation: The lower bound of the solution improves proportionally to the oscillation from $B_{r}$ to $B_{\rho r}$, if $\fint_{B_{r}}u\geq\mu_{r}$. Otherwise the upper bound improves by the same amount. Then we repeat the argument from $B_{\rho r}$ to $B_{\rho^{2}r}$ and so on.
Proof.
Our goal is to show that for some $\rho,\theta\in(0,1)$ we have that $\operatorname*{osc}_{B_{\rho^{k}}}u\leq(1-\theta)^{k}$ holds true for any integer $k\geq 0$, so that Lemma 2.1 and the Remark 2.2 would settle the proof.
Let $\delta\in(0,1)$ as in Lemma 2.6. Assume that $\operatorname*{osc}_{B_{1}}u\leq 1$ and $\|\Delta u\|_{L^{p}(B_{1})}\leq\delta$, otherwise apply the following proof to the function $\bar{u}:=u/\left(\operatorname*{osc}_{B_{1}}u+\delta^{-1}\|\Delta u\|_{L^{p}(B_{1})}\right)$ which does satisfies the required conditions.
Assume by induction that, for $\rho:=\min(1/3,(1-\theta)^{1/(2-n/p)})$ and some $k\geq 1$, we have that $\operatorname*{osc}_{B_{\rho^{k}}}u\leq(1-\theta)^{k}$. The base case $k=1$ is a just the conclusion of Lemma 2.6. Then consider $v(x):=(1-\theta)^{-k}u(\rho^{k}x)$ such that
$\begin{cases}\operatorname*{osc}_{B_{1}}v=(1-\theta)^{-k}\operatorname*{osc}_{B_{\rho^{k}}}u\leq 1,\\ \|\Delta v\|_{L^{p}(B_{1})}=(1-\theta)^{-k}\rho^{k(2-n/p)}\|\Delta u\|_{L^{p}(B_{\rho^{k}})}\leq\delta.\end{cases}$
Then by Lemma 2.6, $\operatorname*{osc}_{B_{\rho}}v\leq\operatorname*{osc}_{B_{1/3}}v\leq(1-\theta)$. Equivalently,
$\operatorname*{osc}_{B_{\rho^{k+1}}}u=(1-\theta)^{k}\operatorname*{osc}_{B_{\rho}}v\leq(1-\theta)^{k+1},$
which now shows that the geometric decay is valid for every $k\geq 1$. ∎
The final proof of Theorem 2.3 now follows from Lemma 2.7 by an idea which is usually referred in the literature as a covering argument. The details are usually omitted as they are routinary and widely known by experts. We give the details in the next demonstration.
Proof of Theorem 2.3.
Let $B_{r}(x_{0})\subseteq\Omega$, $x,y\in B_{r/6}(x_{0})$, and $\rho=|x-y|\in(0,r/3)$. Notice that $B_{r/3}(x)\subseteq B_{r}(x_{0})\subseteq\Omega$. By the previous corollary applied to $v(z):=u((r/3)z+x)$, we know that
$\frac{|u(y)-u(x)|}{|y-x|^{\alpha}}$ $\leq\rho^{-\alpha}\operatorname*{osc}_{B_{\rho}(x)}u$
$=\rho^{-\alpha}\operatorname*{osc}_{B_{3\rho/r}}v$
$\leq Cr^{-\alpha}\left(\operatorname*{osc}_{B_{1}}v+\|\Delta v\|_{L^{p}(B_{1})}\right)$
$\leq Cr^{-\alpha}\left(\operatorname*{osc}_{\Omega}u+r^{2-n/p}\|\Delta u\|_{L^{p}(B_{r/2}(x_{0}))}\right)$
$\leq Cr^{-\alpha}\left(\operatorname*{osc}_{\Omega}u+\|\Delta u\|_{L^{p}(\Omega)}^{(2-n/p)}\right).$
By taking the supremum over $x,y\in B_{r/6}(x_{0})$
$r^{\alpha}[u]_{C^{0,\alpha}(B_{r/6}(x_{0}))}\leq C\left(\operatorname*{osc}_{\Omega}u+\|\Delta u\|_{L^{p}(\Omega)}^{(2-n/p)}\right).$
Now it suffices to show that for some constant $C\geq 1$
$[u]_{C^{0,\alpha}(\Omega)}^{(\alpha)}\leq C\sup_{B_{\rho}(x)\subseteq\Omega}\rho^{\alpha}[u]_{C^{0,\alpha}(B_{\rho/6}(x))}.$
The difference with the original definition for the weighted semi-norm is that in the right-hand side we have the radius $\rho/6$ instead of $\rho/2$.
Given $B_{r}(x_{0})\subseteq\Omega$, consider a covering of $B_{r/2}(x_{0})$ with balls of radius $r/12$, centered at some point in $B_{r/2}(x_{0})$. We can do this covering with $N$ of balls, where $N$ only depends on the dimension. If $B_{r/12}(y)$ is one of these balls we get that $B_{r/2}(y)\subseteq B_{r}(x_{0})\subseteq\Omega$, then
$r^{\alpha}[u]_{C^{0,\alpha}(B_{r/12}(y))}\leq 2^{\alpha}\sup_{B_{\rho}(x)\subseteq\Omega}\rho^{\alpha}[u]_{C^{0,\alpha}(B_{\rho/6}(x))}.$
By sub-additivity of the semi-norm we get that
$r^{\alpha}[u]_{C^{0,\alpha}(B_{r/2}(x_{0}))}\leq 2^{\alpha}N\sup_{B_{\rho}(x)\subseteq\Omega}\rho^{\alpha}[u]_{C^{0,\alpha}(B_{\rho/6}(x))}.$
Then the result follows after taking the supremum over $B_{r}(x_{0})\subseteq\Omega$ on the left-hand side. ∎
Remark 2.8.
Notice how the Hölder estimate obtained from $\Delta u\in L^{p}$ with $p>n/2$ is related to the sequential application of Sobolev’s embedding and Morrey’s embedding, $W^{2,p}\subseteq W^{1,q}\subseteq C^{0,\alpha}$ for $p>n/2$, $q=np/(n-p)_{+}$, and $\alpha=1-n/q$ (if $p\geq n$ we take $q=\infty$ and $\alpha=1$). The surprising observation is that the Laplacian only provides partial information on the Hessian.
Remark 2.9.
For $\alpha>0$, the function $u(x):=|\ln|x||^{-\alpha}$ over $B_{1/2}\subseteq{\mathbb{R}}^{n}$ is not Hölder continuous, however
$\Delta u(x)\sim|\ln|x||^{-\alpha-2}|x|^{-2}\in L^{n/2}(B_{1/2}).$
This shows that the critical exponent $p=n/2$ can not be reached in the Hölder estimate.
Corollary 2.10 (Interior Higher Order Estimates - Laplacian).
Given $p>n/2$ and a non-negative integer $k$, there exist $\alpha\in(0,1)$ and $C\geq 1$, such that for $\Omega\subseteq{\mathbb{R}}^{n}$ open, and $u\in C^{k+2}(\Omega)$ we have that
$[D^{k}u]^{(k+\alpha)}_{C^{0,\alpha}(\Omega)}\leq C\left(\operatorname*{osc}_{\Omega}u+\|D^{k}(\Delta u)\|_{L^{p}(\Omega)}^{(k+2-n/p)}\right).$
Let us present the proof for $k=1$, as the general case follows from a similar reasoning.
Proof.
For any directional derivative $v:=\partial_{e}u$, with respect to a unit direction $e\in\partial B_{1}$, we have that $\Delta v=\partial_{e}(\Delta u)$ in $\Omega$. By applying the proof of Theorem 2.3 to $v$ we get that for any $B_{r}(x_{0})\subseteq\Omega$
$r^{\alpha}[v]_{C^{0,\alpha}(B_{r/6}(x_{0}))}$ $\leq C\left(\operatorname*{osc}_{B_{r/2}(x_{0})}v+r^{2-n/p}\|\partial_{e}(\Delta u)\|_{L^{p}(B_{r/2}(x_{0}))}\right)$
$\leq C\left(\|Du\|_{L^{\infty}(B_{r/2}(x_{0}))}+r^{2-n/p}\|D(\Delta u)\|_{L^{p}(B_{r/2}(x_{0}))}\right).$
By taking the supremum over $e\in\partial B_{1}$ on the left-hand side
(2.1) $r^{\alpha}[Du]_{C^{0,\alpha}(B_{r/6}(x_{0}))}\leq C\left(\|Du\|_{L^{\infty}(B_{r/2}(x_{0}))}+r^{2-n/p}\|D(\Delta u)\|_{L^{p}(B_{r/2}(x_{0}))}\right).$
Now we use the an interpolation argument to split the term $\|Du\|_{L^{\infty}(B_{r/2}(x_{0}))}$ between the oscillation of $u$ and a sufficiently small multiple of the left-hand side: Let $\delta\in(0,1/4)$ a small constant to be fixed, and $x_{1}\in B_{r/2}(x_{0})$ arbitrary. As a first step we will find a point $x_{2}\in B_{\delta r}(x_{1})$ where $|Du(x_{2})|\leq C\operatorname*{osc}_{\Omega}u/(\delta r)$. Notice that in dimension $n=1$ this follows from Rolle’s theorem.
Consider the paraboloid
$\varphi(x):=\min_{\overline{B_{\delta r}(x_{1})}}u+\frac{\operatorname*{osc}_{\Omega}u}{\delta^{2}r^{2}}\left(\delta^{2}r^{2}-|x-x_{1}|^{2}\right).$
Let us see that there exists some $x_{2}\in\operatorname*{argmin}_{B_{\delta r}(x_{1})}(u-\varphi)$, see Figure 5.
Notice that $m:=\min_{\overline{B_{\delta r}(x_{1})}}(u-\varphi)$ must be non-negative because $\varphi(x_{1})\geq u(x_{1})$. If $m=0$ we just take $x_{2}=x_{1}$. Otherwise, if $m<0$, then $K:=\{u\leq\varphi+m\}\cap B_{\delta r}(x_{1})$ is a non-empty compact set and by Weierstrass theorem we get that
$\emptyset\neq\operatorname*{argmin}_{K}(u-\varphi)=\operatorname*{argmin}_{B_{\delta r}(x_{1})}(u-\varphi)\subseteq B_{\delta r}(x_{1}).$
By applying the first derivative test at $x_{2}\in\operatorname*{argmin}_{B_{\delta r}(x_{1})}(u-\varphi)$ we obtain that
$|Du(x_{2})|\leq 2\delta^{-1}r^{-1}\operatorname*{osc}_{\Omega}u.$
Then
$|Du(x_{1})|\leq|Du(x_{2})|+|Du(x_{1})-Du(x_{2})|\leq Cr^{-1}\operatorname*{osc}_{\Omega}u+\delta^{\alpha}r^{\alpha}[Du]_{C^{0,\alpha}(B_{3r/4}(x_{0}))}.$
By taking the supremum over $x_{1}\in B_{r/2}(x_{0})$ we get
$\|Du\|_{L^{\infty}(B_{r/2}(x_{0}))}\leq Cr^{-1}\operatorname*{osc}_{\Omega}u+\delta^{\alpha}r^{\alpha}[Du]_{C^{0,\alpha}(B_{3r/4}(x_{0}))}.$
Substituting this interpolation in (2.1), and multiplying by $r$, we get that
$r^{1+\alpha}[Du]_{C^{0,\alpha}(B_{r/6}(x_{0}))}\leq C\left(\operatorname*{osc}_{\Omega}u+\delta^{\alpha}r^{1+\alpha}[Du]_{C^{0,\alpha}(B_{3r/4}(x_{0}))}+\|D(\Delta u)\|^{(3-n/p)}_{L^{p}(\Omega)}\right).$
Figure 5. The following construction could be considered a multidimensional version of Rolle’s theorem. The paraboloid $\varphi$ satisfies that $\varphi(x_{1})\geq u(x_{1})$ and $\varphi\leq u$ on $\Omega\setminus B_{\delta r}(x_{1})$. This means that by continuity and compactness, the graph of $u$ must cross the graph of $\varphi$ in $B_{\delta r}(x_{1})$. Then a downwards translation of $\varphi$ becomes tangent with $u$ at some point $x_{2}\in B_{\delta r}(x_{1})$. At such contact point the gradient of $u$ is comparable with $\operatorname*{osc}_{\Omega}u/(\delta r)$.
To conclude let us point out that the argument at the end of the proof of Theorem 2.3 shows that
$[Du]^{(1+\alpha)}_{C^{0,\alpha}(\Omega)}\leq C\sup_{B_{r}(x_{0})\subseteq\Omega}r^{1+\alpha}[Du]_{C^{0,\alpha}(B_{r/6}(x_{0}))}.$
In a similar way we can also show that
$\sup_{B_{r}(x_{0})\subseteq\Omega}r^{1+\alpha}[Du]_{C^{0,\alpha}(B_{3r/4}(x_{0}))}\leq C[Du]^{(1+\alpha)}_{C^{0,\alpha}(\Omega)}.$
Hence, by finally choosing $\delta$ sufficiently small we get to absorb the term $C\delta^{\alpha}[Du]^{(1+\alpha)}_{C^{0,\alpha}(\Omega)}$ on the left-hand side of the estimate. ∎
{exercise}
[Morrey’s Inequality. The identities in this problem could be considered as analogues of the mean value property for the first order operator $|Du|$.]
(Dimension 1) Show that for $p>1$, $u\in C^{1}((0,1))$, and $\alpha=1-1/p$
$[u]_{C^{0,\alpha}(0,1)}\leq\|u^{\prime}\|_{L^{p}(0,1)}.$
(General dimension, integral identity) Prove that for some constant $C\geq 1$ it holds that for any $r\in(0,1)$ and $u\in C^{1}(B_{1})$
[…]
(General dimension, Hölder estimate) Given $p>n$, there exists some constant $C\geq 1$ such that for $u\in C^{1}(B_{1})$
$\sup_{r\in(0,1/2)}r^{-(1-n/p)}\operatorname*{osc}_{B_{r}}u\leq C\|Du\|_{L^{p}(B_{1})}.$
{exercise}
Show that for a non-negative harmonic function $u\in C^{2}(B_{1})$ and $r\in(0,1/3)$,
$\sup_{B_{r}}u\leq\left(\frac{1-r}{1-3r}\right)^{n}\inf_{B_{r}}u.$
Use this previous inequality to show a diminish of oscillation and interior Hölder’s estimate for harmonic functions considering as alternatives whether $u(0)$ is smaller or greater than $(\inf_{B_{1/4}}u+\sup_{B_{1/4}}u)/2$.
In these notes we focus on a particular proof of the Harnack inequality. Other remarkable strategies that can also be extended to non-divergence uniformly elliptic equations include: The proof by Serrin [Ser56] using barriers in two dimensions, showcased in [GT83, Theorem 3.10]. The one by Li and Yau [LY86], using the Bernstein method and presented in the simplest scenario in [HL11, Lemma 1.32].
{exercise}
[Concave/Convex Functions. Many properties for sub/super-harmonic functions have easier counterparts for convex/concave functions. Try to find easy proof for the following facts, without using that convex/concave functions are sub/super-harmonic.]
(Harnack) For $u\in C(B_{1})$ non-negative and concave, and $r\in(0,1)$
$\sup_{B_{1}}u\leq\frac{2}{1-r}\inf_{B_{r}}u.$
(Local Maximum Principle) For every $u\in C(B_{1})\cap L^{1}(B_{1})$ non-negative and convex, and $r\in(0,1)$
$\sup_{B_{r}}u\leq\frac{2^{n}}{|B_{1}|(1-r)^{n}}\|u\|_{L^{1}(B_{1}\setminus B_{r})}.$
{exercise}
[Weyl’s Lemma. From Lemma 2.4 we get that a harmonic function $u$ must satisfy the mean value property
$u(x_{0})=\fint_{B_{r}(x_{0})}u.$
On the other hand, this identity is rigid enough to imply smoothness and harmonicity. We say that $u\in L^{1}(\Omega)$ is harmonic in a weak sense if and only if the mean value property holds for every ball $B_{r}(x_{0})\subseteq\Omega$.]
Let $\eta\in C^{\infty}_{c}(B_{1})$ non-negative, radially symmetric, and such that $\int_{B_{1}}\eta=1$; and let $\eta_{\varepsilon}(x):=\varepsilon^{-n}\eta(\varepsilon^{-1}x)$. Show that if $u\in L^{1}(\Omega)$ is weakly harmonic, then $u=\eta_{\varepsilon}\ast u$ in $\Omega_{\varepsilon}:=\{x\in\Omega\,:\,B_{\varepsilon}(x)\subseteq\Omega\}$. This fact can be used to show that $u$ is locally smooth.
Show that there exists a constant $C\geq 1$, depending only on the dimension, such that for every $u\in C^{2}(B_{\varepsilon})$
$\lim_{\rho\to 0^{+}}\rho^{-2}\fint_{B_{\rho}}(u-u(0))=C\Delta u(0).$
This finally shows that if $u$ is weakly harmonic, then it must also be harmonic in the classical sense.
{exercise}
Show that for every $p>n/2$ and $\eta\in(0,1)$ there exist $\alpha\in(0,1)$ and $C\geq 1$ such that the following holds: Let $\Omega\subseteq B_{1}$ such that
$\inf_{r\in(0,1)}\frac{|({\mathbb{R}}^{n}\setminus\Omega)\cap B_{r}|}{|B_{r}|}\geq\eta.$
Then for any $u\in C^{2}(\Omega)\cap C(B_{1})$ with $u=0$ in $B_{1}\setminus\Omega$ and $\Delta u\in L^{p}(\Omega)$, we have that
$\sup_{r\in(0,1)}r^{-\alpha}\operatorname*{osc}_{B_{r}}u\leq C\left(\operatorname*{osc}_{B_{1}}u+\|\Delta u\|_{L^{p}(\Omega)}\right).$
{exercise}
Show that for every $p>n/2$ and $\varepsilon>0$, there exist $\theta\in(0,1)$ and $M\geq 1$ such that the following holds: For $u\in C^{2}(B_{1})$ with $\|u_{+}\|_{L^{\varepsilon}(B_{1})}\leq 1$, and $\|(\Delta u)_{-}\|_{L^{p}(B_{1})}\leq 1$, we have that
$u(0)\geq M\qquad\Rightarrow\qquad\sup_{B_{1}}u\geq(1+\theta)u(0).$
Show that for every $p>n/2$ and $\varepsilon>0$, there exists some $C\geq 1$ such that the following maximum principle holds for $u\in C^{2}(B_{1})$
$\max_{\overline{B_{1/2}}}u_{+}\leq C\left(\|u_{+}\|_{L^{\varepsilon}(B_{1})}+\|(\Delta u)_{-}\|_{L^{p}(B_{1})}\right).$
This implies that the interior Hölder estimate can be strengthened to [^4] For $\varepsilon\in(0,1)$, $\|\cdot\|_{L^{\varepsilon}}$ is not a norm because it is not sub-additive, however it is sub-additive up to a multiplicative constant.
$[u]_{C^{0,\alpha}(\Omega)}^{(\alpha)}\leq C\left(\|u\|_{L^{\varepsilon}(\Omega)}^{(-n/\varepsilon)}+\|\Delta u\|_{L^{p}(\Omega)}^{(2-n/p)}\right).$
{exercise}
[Heat Equation. Let $H_{t}(x):=(4\pi t)^{-n/2}e^{-|x|^{2}/(4t)}$ be the heat kernel. For these problems it may be convenient to know that $u=u(x,t)\colon{\mathbb{R}}^{n}\times(-\infty,0]\to{\mathbb{R}}$ defined from $f=f_{t}(x)\in C_{c}({\mathbb{R}}^{n}\times(-\infty,0])$ by the global representation formula
[…]
is the unique bounded solution of the heat equation $\partial_{t}u-\Delta u=f$ in ${\mathbb{R}}^{n}\times(0,-\infty]$, with $u(x,-T)=0$ for any $T>0$ such that $\operatorname*{spt}f\subseteq{\mathbb{R}}^{n}\times(-T,0]$.]
(Mean Value Property) Show that there exist some non-negative kernels $K_{1},K_{2}\in C_{c}(B_{1}\times(-1,0])$, such that for any solution $u\in(C^{2}_{x}\cap C^{1}_{t})(B_{1}\times(-1,0])$ of the heat equation $\partial_{t}u-\Delta u=f$ in $B_{1}\times(0,1]$ it holds that
$u(0,0)=\iint_{B_{1}\times(-1,0]}(uK_{1}+fK_{2}).$
(Weak Harnack) Show that for every $p,q\in[1,\infty]$ such that $2>n/p+2/q$, there exist a radius $\rho\in(0,1/4)$ and a constant $C\geq 1$ such that for any non-negative solution $u\in(C^{2}_{x}\cap C^{1}_{t})(B_{1}\times(-1,0])$ of $\partial_{t}u-\Delta u=f$ in $B_{1}\times(-1,0]$ it holds that
$\iint_{Q_{-}}u\leq C\left(\inf_{Q_{+}}u+\|f_{-}\|_{L^{q}_{t}((-1,0]\to L^{p}_{x}(B_{1}))}\right),$
where $Q_{-}:=B_{\rho}\times(-4\rho,-3\rho]$ and $Q_{+}:=B_{\rho}\times(-\rho,0]$. Similarly to the stationary problem, this estimate yields an interior Hölder estimate for the solution of a heat equation of the form
$\sup_{r\in(0,1)}r^{-\alpha}\operatorname*{osc}_{B_{r}\times(-r^{2},0]}u\leq C\left(\operatorname*{osc}_{B_{1}\times(-1,0]}u+\|f\|_{L^{q}_{t}((-1,0]\to L^{p}_{x}(B_{1}))}\right).$
3. Uniformly Elliptic Equations
3.1. Preliminary Observations
Let $\Omega\subseteq{\mathbb{R}}^{n}$ be open and $F=F(M,p,z,x)\in C({\mathbb{R}}^{n\times n}_{\text{sym}}\times{\mathbb{R}}^{n}\times{\mathbb{R}}\times\Omega)$ be a real valued function with which we specify a second-order non-linear PDE of the form
$F(D^{2}u(x),Du(x),u(x),x)=0\text{ for }x\in\Omega.$
To simplify the notation we may just write the expression as
$F(D^{2}u,Du,u,x)=0\text{ in $\Omega$}.$
In analogy with the definition of super-harmonic functions, we call super-solutions those functions that satisfy the inequality $F(D^{2}u,Du,u,x)\leq 0$. Similarly, sub-solutions are those that satisfy the inequality $F(D^{2}u,Du,u,x)\geq 0$. A mnemotechnic rule to recall the signs is that super-solutions tend to be concave, the prefix super could be associated with the fact that their graph tend to curve upwards, however their Hessians then tend to be negative which explains the direction in the inequality.
3.1.1. Linearizations
A common strategy to analyze non-linear equation consists of linearizing the problem at hand, leading to a problem of the form [^5] We assume the following inner product in ${\mathbb{R}}^{n\times n}$ $A:B:=\operatorname{tr}(A^{T}B)=\sum_{i,j=1}^{n}a_{ij}b_{ij}.$
(3.2) $Lv:=A:D^{2}v+b\cdot Dv+cv=f.$
Occasionally we will use $L=A:M+b\cdot p+cz$ to refer to a linear operator as above.
The coefficients given by the diffusion matrix $A$, the drift vector $b$, the scalar $c$, and the forcing term $f$, are usually functions that depend on the spatial variable $x$. In such instances, we describe the operator as having variable coefficients. Notice that the simplest case $A=I$, $b=0$ and $c=0$ makes $L=\Delta$. We will consider three ways to obtain this linearization, depending on whether we look for estimates on $u$, the gradient of $u$, or the Hessian of $u$.
In the following we assume that $F\in C^{1}({\mathbb{R}}^{n\times n}_{\text{sym}}\times{\mathbb{R}}^{n}\times{\mathbb{R}}\times\Omega)$ is differentiable.
Linearization for the Original Function: The first way follows from the fundamental theorem of calculus. By integrating the derivative $(d/dt)F(tD^{2}u,tDu,tu,x)$ from $t=0$ to $t=1$,
[…]
$=\sum_{i,j=1}^{n}\left(\int_{0}^{1}\partial_{m_{ij}}Fdt\right)\partial_{ij}u+\sum_{i=1}^{n}\left(\int_{0}^{1}\partial_{p_{i}}Fdt\right)\partial_{i}u+\left(\int_{0}^{1}\partial_{z}Fdt\right)u.$
We obtain in this way that $v_{0}=u$ satisfies the linear equation (3.2), with coefficients given by
[…]
$c(x)$ $:=\int_{0}^{1}\partial_{z}F(tD^{2}u(x),tDu(x),tu(x),x)dt,$
$f(x)$ $:=-F(0,0,0,x).$
The functions $D_{M}F=(\partial_{m_{ij}}F)\in{\mathbb{R}}^{n\times n}_{\text{sym}}$, $D_{p}F=(\partial_{p_{1}}F,\ldots,\partial_{p_{n}}F)\in{\mathbb{R}}^{n}$, and $\partial_{z}F\in{\mathbb{R}}$ denote the partial derivatives of $F$ with respect to the Hessian variable $M=(m_{ij})\in{\mathbb{R}}^{n\times n}_{\text{sym}}$, the gradient variable $p=(p_{1},\cdots,p_{n})\in{\mathbb{R}}^{n}$, and $z\in{\mathbb{R}}$, respectively.
Linearization for a First Derivative A second way to obtain a similar linearization consists of taking a derivative of the equation in some unit direction $e\in\partial B_{1}$. This leads to the following linear equation for $v_{1}:=\partial_{e}u$
$D_{M}F:D^{2}v_{1}+D_{p}F\cdot Dv_{1}+\partial_{z}Fv_{1}=-e\cdot D_{x}F.$
In other words, $v_{1}$ satisfies $Lv_{1}=f$ for the following coefficients
$A(x)$ $:=D_{M}F(D^{2}u(x),Du(x),u(x),x),$
$b(x)$ $:=D_{p}F(D^{2}u(x),Du(x),u(x),x),$
$c(x)$ $:=\partial_{z}F(D^{2}u(x),Du(x),u(x),x)\partial_{e}u(x),$
$f(x)$ $:=-e\cdot D_{x}F(D^{2}u(x),Du(x),u(x),x).$
Linearization for a Second Derivative We can also obtain a linear equation for a pure second derivative $v_{2}=\partial_{e}^{2}u$ by differentiating the equation for $u$ twice, or equivalently the equation for $v_{1}=\partial_{e}u$ once. Keep in mind that information about mixed derivatives can be recovered from the polarization identity
$2\partial_{e^{\prime}}\partial_{e}=\partial_{e+e^{\prime}}^{2}-\partial_{e}^{2}-\partial_{e^{\prime}}^{2}.$
Let us focus on the effect of the derivative on the leading order term of the equation for $v_{1}$
$\partial_{e}\left(\sum_{i,j=1}^{n}\partial_{m_{ij}}F\partial_{ij}v_{1}\right)=\sum_{i,j=1}^{n}\partial_{m_{ij}}F\partial_{ij}v_{2}+\sum_{i,j,k,l=1}^{n}(\partial_{m_{kl}}\partial_{m_{ij}}F)\partial_{ij}v_{1}\partial_{kl}v_{1}.$
Notice that the second term above have a sign if we assume that $F$ is concave or convex in the Hessian variable.
We can repeat the same procedure to obtain linear equations for higher order derivatives of $u$. However, once we have established $C^{3}$ estimates for the solution, higher order estimates become readily available thanks to the Schauder theory, not covered in this note and for which we recommend [FRRO22, Chapter 2]. This technique is known as elliptic bootstrapping, before describing it we must understand the most important hypotheses, ellipticity and uniform ellipticity.
3.1.2. Ellipticity and Uniform Ellipticity
In the previous linearizations the coefficients depend on the solution $u$ as well. However, under suitable hypotheses on the derivatives of $F$ and $u$, we can overlook this dependence and understand the equation in a general sense. For instance, it may be the case that the coefficients are all measurable and bounded.
A common assumption in many applications is the ellipticity of the operator. This appears as a monotonicity property given by [^6] We assume the following ordering in ${\mathbb{R}}^{n\times n}_{\text{sym}}$: $M\leq N$ if and only if $x\cdot Mx\leq x\cdot Nx$ for every $x\in{\mathbb{R}}^{n}$. Equivalently, the eigenvalues of $(N-M)$ are non-negative. $D_{M}F\geq 0$. Uniform ellipticity requires instead a quantitative form of monotonicity with respect to the Hessian variable
(3.3) $\lambda I\leq D_{M}F\leq\Lambda I,$
for some parameters $[\lambda,\Lambda]\subseteq(0,\infty)$.
We can also extend the notions of ellipticity and uniform ellipticity to non-differentiable operators. Ellipticity means that $F$ is monotone in the Hessian variable, meaning that for $M,N\in{\mathbb{R}}^{n\times n}_{\text{sym}}$ and $(p,z,x)\in{\mathbb{R}}^{n}\times{\mathbb{R}}\times\Omega$ we have that
$M\leq N\qquad\Rightarrow\qquad F(M,p,z,x)\leq F(N,p,z,x).$
The main consequence of this property is the comparison principle.
Figure 6. The graph of $u$ touches the graph of $v$ from below. Over the contact point $x_{0}$ we obtain that $u(x_{0})=v(x_{0})$, $Du(x_{0})=Dv(x_{0})$ (first derivative test), and $D^{2}u(x_{0})\leq D^{2}v(x_{0})$ (second derivative test).
The Comparison Principle: Given two functions $u,v\colon B_{r}(x_{0})\to{\mathbb{R}}$, we say that $u$ touches $v$ from below at $x_{0}\in\Omega$ if $u\leq v$ in $B_{r}(x_{0})$, and $u(x_{0})=v(x_{0})$ (Figure 6). Whenever the functions are differentiable at $x_{0}$, the first derivative test implies that their gradients at the contact point are the same. If additionally we assume that the functions have second order derivatives at $x_{0}$, then the second derivative test says that their Hessians are ordered, $D^{2}u(x_{0})\leq D^{2}v(x_{0})$. For $F$ elliptic it then must happen that
$F(D^{2}u(x_{0}),Du(x_{0}),u(x_{0}),x_{0})\leq F(D^{2}v(x_{0}),Dv(x_{0}),v(x_{0}),x_{0}).$
The comparison principle motivates the notion of viscosity solutions, a theory of weak solutions developed during the eighties for first and second-order problems and widely used nowadays. Even though they will not be used in these notes, it is useful to describe the general construction and keep it in mind for the coming section about the contact set for a whole family of test functions (Section 3.2.3).
In the case of viscosity super-solutions, this notion declares that $u\in C(\Omega)$ satisfies
$F(D^{2}u,Du,u,x)\leq 0\text{ in the viscosity sense $\Omega$},$
by checking that for any $B_{r}(x_{0})\subseteq\Omega$, and any test function $\varphi\in C^{2}(B_{r}(x_{0}))$ touching $u$ from below at $x_{0}$, it holds that at the contact point
$F(D^{2}\varphi(x_{0}),D\varphi(x_{0}),\varphi(x_{0}),x_{0})\leq 0.$
The ellipticity of $F$ means that this criteria is automatically satisfied if $u\in C^{2}(\Omega)$ is a super-solution in the classical sense.
Viscosity sub-solutions require instead a test function $\varphi$ touching $u$ from above at some $x_{0}\in\Omega$, and checking that $F(D^{2}\varphi(x_{0}),D\varphi(x_{0}),\varphi(x_{0}),x_{0})\geq 0$. The equality in the viscosity sense holds when the function is simultaneously a viscosity super and a sub-solution. Further details about this theory can be consulted in [CIL92, CC95, Cal18].
In the literature, the comparison principle is presented as result that allows to compare a sub and super-solution ordered over the boundary of a bounded domain.
To be precise, let $\Omega\subseteq{\mathbb{R}}^{n}$ be open and bounded, and $F\in C({\mathbb{R}}^{n\times n}_{\text{sym}}\times{\mathbb{R}}^{n}\times{\mathbb{R}}\times\Omega)$ elliptic and non-increasing in the $z$ variable, this last property is known as properness. Given $u,v\in C^{2}(\Omega)\cap C(\overline{\Omega})$ we get that
$\begin{cases}F(D^{2}u,Du,u,x)>F(D^{2}v,Dv,v,x)\text{ in }\Omega,\\ u\leq v\text{ in }\partial\Omega,\end{cases}\qquad\Rightarrow\qquad u<v\text{ in }\Omega.$
Figure 7. If the graphs of $u$ and $v$ cross or touch each other, we get that some downwards translation $u-M$ touches $v$ from below at some $x_{0}\in\Omega$. Given that $M\geq 0$ and $F$ is elliptic and proper, this leads to the contradiction $F(D^{2}u(x_{0}),Du(x_{0}),u(x_{0}),x_{0})\leq F(D^{2}u(x_{0}),Du(x_{0}),u(x_{0})-M,x_{0})\leq F(D^{2}v(x_{0}),Dv(x_{0}),v(x_{0}),x_{0})$.
In contrast to the direct reasoning for classical sub and super-solutions (Figure 7), the comparison principle for viscosity solutions is a highly non-trivial result that took several years to be fully established.
The Extremal Pucci Operators: Let $F\in C^{1}({\mathbb{R}}^{n\times n}_{\text{sym}}\times{\mathbb{R}}^{n}\times{\mathbb{R}}\times\Omega)$ be uniformly elliptic. By the fundamental theorem of calculus
[…]
Using that $\lambda I\leq D_{M}F\leq\Lambda I$, we deduce the identity
(3.4) $\min_{\lambda I\leq A\leq\Lambda I}A:(M-N)\leq F(M,p,z,x)-F(N,p,z,x)\leq\max_{\lambda I\leq A\leq\Lambda I}A:(M-N).$
This property is actually equivalent to $\lambda I\leq D_{M}F\leq\Lambda I$ and does not requires differentiablity on $F$. Therefore, it can be used to extend the notion of uniform ellipticity when $F$ is not differentiable.
The operators that appear on the left and right-hand side of (3.4) are the extremal Pucci operators and can be computed as the following combinations of the eigenvalues
$\mathcal{P}^{-}_{\lambda,\Lambda}(M)$ $:=\min_{\lambda I\leq A\leq\Lambda I}A:M=\sum_{e\in\operatorname{eig}(M)}\lambda e_{+}-\Lambda e_{-},$
$\mathcal{P}^{+}_{\lambda,\Lambda}(M)$ $:=\max_{\lambda I\leq A\leq\Lambda I}A:M=\sum_{\lambda\in\operatorname{eig}(M)}\Lambda e_{+}-\lambda e_{-}.$
Remark 3.1.
The Pucci operators provide an indirect way to address equations of the form $A:D^{2}u=f$ without a explicit characterization of the coefficients given by the matrix valued function $A$, besides uniform ellipticity. Indeed, for $f\in C(\Omega)$ we have that $u\in C^{2}(\Omega)$ satisfies
$\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u)\leq f\leq\mathcal{P}^{+}_{\lambda,\Lambda}(D^{2}u)\text{ in $\Omega$}$
[…]
The homogeneous equation $\mathcal{P}_{\lambda,\Lambda}^{-}(D^{2}u)\leq 0$ states that the positive eigenvalues of the Hessian are controlled by the negative ones. Meanwhile the equation $\mathcal{P}_{\lambda,\Lambda}^{+}(D^{2}u)\geq 0$ states the that the negative eigenvalues are controlled by the positive ones. (Figure 8)
Figure 8. For the homogeneous equation $\mathcal{P}_{\lambda,\Lambda}^{-}(D^{2}u)\leq 0\leq\mathcal{P}_{\lambda,\Lambda}^{+}(D^{2}u)$ we get that the graph of $u$ has a saddle at every point. Moreover, the sum of the positive eigenvalues of the Hessian and the sum of the negative eigenvalues of the Hessian are comparable: $(\lambda/\Lambda)\sum_{e\in\operatorname{eig}(D^{2}u)}e_{-}\leq\sum_{e\in\operatorname{eig}(D^{2}u)}e_{+}\leq(\Lambda/\lambda)\sum_{e\in\operatorname{eig}(D^{2}u)}e_{-}$.
Remark 3.2.
For a uniformly elliptic operator $F=F(M,x)$, not necessarily differentiable, we can “linearize” the equation $F(D^{2}u,x)=0$ without using the fundamental theorem of calculus. Indeed, if we let $f(x):=-F(0,x)$, then any solution of $F(D^{2}u,x)=0$ satisfies
$\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u)\leq\underbrace{F(D^{2}u,x)-F(0,x)}_{=f}\leq\mathcal{P}^{+}_{\lambda,\Lambda}(D^{2}u).$
If $F=F(M)$ is also independent of $x$, then we can find a linearization for a function approximating a directional derivative $v_{1}=\partial_{e}v$ in the following way:
$v_{1,h}:=\frac{u_{h,e}-u}{h},\qquad u_{h,e}(x):=u(x+he),\qquad h>0,\qquad e\in\partial B_{1}.$
We notice that if $u$ is a solution of $F(D^{2}u)=0$ in $\Omega$, then also the translation $u_{h,e}$ is a solution of the same equation in $\Omega-he$. On the intersection we get that
$\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}v_{1,h})\leq\underbrace{\frac{F(D^{2}u_{h,e})-F(D^{2}u)}{h}}_{=0}\leq\mathcal{P}^{+}_{\lambda,\Lambda}(D^{2}v_{1,h}).$
If we assume that $u\in C^{3}(\Omega)$, then as we send $h\to 0^{+}$ we get that $\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}v_{1})\leq 0\leq\mathcal{P}^{+}_{\lambda,\Lambda}(D^{2}v_{1})$.
Finally, if $F=F(M)$ is uniformly elliptic and convex we get a super-solution inequality for
$v_{2,h}:=\frac{u_{h,e}+u_{-h,e}-2u}{h^{2}}\sim v_{2}=\partial_{e}^{2}u.$
By uniform ellipticity
$\frac{h^{2}}{2}\mathcal{P}^{-}_{\lambda,\Lambda}\left(D^{2}v_{2,h}\right)$ $=\mathcal{P}^{-}_{\lambda,\Lambda}\left(\frac{D^{2}u_{h,e}+D^{2}u_{-h,e}}{2}-D^{2}u\right)$
$\leq F\left(\frac{D^{2}u_{h,e}+D^{2}u_{-h,e}}{2}\right)-F(D^{2}u)$
$=F\left(\frac{D^{2}u_{h,e}+D^{2}u_{-h,e}}{2}\right).$
By convexity and translation invariance
$F\left(\frac{D^{2}u_{h,e}+D^{2}u_{-h,e}}{2}\right)\leq\frac{F(D^{2}u_{h,e})+F(D^{2}u_{-h,e})}{2}=0.$
We conclude in this way that if $u\in C^{4}(\Omega)$, then $\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}v_{2})\leq 0$.
Let us complete this discussion by illustrating the bootstrapping technique for higher order estimates. Assume that $u$ is a $C^{3}$-regular solution of the uniformly elliptic equation $F(D^{2}u)=0$ with $F$ smooth. Then, any directional derivative $v_{1}=\partial_{e}u$ satisfies a linear uniformly elliptic equation with Lipschitz coefficients. Schauder estimates ([FRRO22, Chapter 2] or [GT83, Chapter 6]) then imply that $v_{1}$ is actually $C^{2,1}$-regular and then $u$ must be $C^{3,1}$-regular. This now implies that the coefficients of the equation satisfied by $v_{1}$ are now even better, $C^{1,1}$-regular, which then imply that the solution $v_{1}$ must be $C^{3,1}$-regular, and $u$ is now promoted to be $C^{4,1}$-regular. Clearly this procedure continues all the way to proving that $u$ is actually smooth.
3.1.3. Symmetries
In our analysis for the Laplacian we used some fundamental symmetries of the operator. These appeared when we proved an estimate in a given configuration and then an analogous result was to a translated and rescaled configuration. For instance, the mean value property (Lemma 2.4) was proved in the unit ball, and then observed that a similar estimate was also valid in arbitrary balls by a change of variable argument.
Let us give a few examples of the fundamental symmetries considered for second-order operators of the form $u\mapsto F(D^{2}u,Du,u,x)$. They have also been implicitly used in the previous considerations about linearizations.
Translation invariance: Holds whenever $F=F(M,p,z)$ is independent of the $x$ variable. In this case, we get that if $f(x):=F(D^{2}u(x),Du(x),u(x))$ and $\bar{u}(x):=u(x+x_{0})$ is a translation of $u$ for some $x_{0}\in{\mathbb{R}}^{n}$, then we can compute $F(D^{2}\bar{u}(x),D\bar{u}(x),\bar{u}(x))=f(x+x_{0})$, with the same translation.
We also consider vertical translations of the graph of $u$ given by $\bar{u}(x):=u(x)+\theta$, for some $\theta\in{\mathbb{R}}$. In this case $F=F(M,p,x)$, independent of the $z$ variable, satisfies that $F(D^{2}\bar{u}(x),D\bar{u}(x),x)=F(D^{2}u(x),Du(x),x)$. Whenever $F$ is proper, i.e. non-increasing in $z$, then we have that for $\theta\geq 0$, it holds that $F(D^{2}\bar{u}(x),D\bar{u}(x),\bar{u}(x),x)\leq F(D^{2}u(x),Du(x),u(x),x)$. This means that upwards translations preserve super-solutions of proper equations. Similarly, downwards translations preserve sub-solutions.
Homogeneity: Holds if for any constant $c\geq 0$, we have that $F(cM,cp,cz,x)=cF(M,p,z,x)$. In this case, we get that if $f(x):=F(D^{2}u(x),Du(x),u(x))$ and $\bar{u}(x):=cu(x)$ is a vertical scaling of $u$, then we can compute $F(D^{2}\bar{u}(x),D\bar{u}(x),\bar{u}(x))=cf(x)$, with the same scaling.
Dilation invariance: Let $\alpha\in(0,1]$. The $\alpha$-Hölder continuity of a function $u\colon B_{1}\to{\mathbb{R}}$ at the origin can be characterized by the scaling transformations given by $u_{r}(x):=r^{-\alpha}u(rx)$, for $r>0$. We are actually interested on the dilations that arise when $r$ belongs to $(0,1)$, so that the graph of $u_{r}$ consists on dilating the graph of $u$ horizontally with a factor $r^{-1}$, and vertically with a factor $r^{-\alpha}$ (Figure 9).
Figure 9. For $r\in(0,1)$, the scaling $u_{r}(x)=r^{-\alpha}u(rx)$ zooms-in the graph of $u$ by identifying the values of $u$ over the ball of radius $r$, to the values of $u_{r}$ over the ball of radius 1. In the figure above we illustrate the case when $\alpha=1$, which corresponds to a similar vertical stretching of the graph.
[…]
$F(r^{\alpha-2}M,r^{\alpha-1}p,r^{\alpha}z,rx)=r^{\beta}F(M,p,z,x).$
In this case, if we let $f(x):=F(D^{2}u(x),Du(x),u(x),x)$ and $u_{r}(x)=r^{-\alpha}u(rx)$ a dilation of $u$, then we are able to compute $F(D^{2}u_{r}(x),Du_{r}(x),u_{r}(x),x)=r^{-\beta}f(rx)$ also in terms of a scaling.
For instance, the Laplacian is invariant by dilations for any pair of exponents $\alpha,\beta\in{\mathbb{R}}$ such that $\beta=\alpha-2$. This is the main consideration behind the change of variable arguments in the diminish of oscillation strategy. Notice as well that any pure second-order linear operator with constant coefficients, as well as the Pucci extremal operators, are invariant by dilations, once again for $\beta=\alpha-2$.
It is also useful to consider the notion of a whole family of operators preserved by some given symmetries. A family of operators =$\mathcal{F}=\{F\}$ is invariant by dilations and in terms of the exponents $\alpha,\beta\in{\mathbb{R}}$, if for any $F\in\mathcal{F}$ and $r\in(0,1)$, there exists $G\in\mathcal{F}$, such that $F(r^{\alpha-2}M,r^{\alpha-1}p,r^{\alpha}z,rx)=r^{\beta}G(M,p,z,x)$.
This previous notion is useful to characterize a family of linear operators $L=A:M+b\cdot p+cz$ with the uniform ellipticity hypothesis $\lambda I\leq A\leq\Lambda I$, and bounded coefficients $\|b\|_{L^{\infty}({\mathbb{R}}^{n})},\|c\|_{L^{\infty}({\mathbb{R}}^{n})}\leq\Lambda$. In general, each one of these operators is not invariant by dilations, however the whole family is. Moreover, the same can be said if instead of the uniform bounds we impose $\|b\|_{L^{q_{1}}({\mathbb{R}}^{n})},\|c\|_{L^{q_{0}}({\mathbb{R}}^{n})}\leq\Lambda$, for $q_{1}\geq n$ and $q_{0}\geq n/2$.
All these symmetries will be combined in order to simplify some arguments in our proofs. The idea starts by showing a result in a normalized domain (usually the unit ball), with some normalized hypothesis (for example, oscillation bounded by one, or a sufficiently small forcing term for the equation). Then the hypotheses and conclusions can be extended to other configurations by applying a combination of translations and rescalings.
3.2. Interior Hölder Estimates for Uniformly Elliptic Equations
Here is the main result of this section and the fundamental estimate of the Krylov-Safonov regularity theory.
Theorem 3.3 (Interior Hölder Estimate - Uniformly Elliptic).
[…]
$[u]^{(\alpha)}_{C^{0,\alpha}(\Omega)}\leq C\left(\operatorname*{osc}_{\Omega}u+\|f\|^{(1)}_{L^{n}(\Omega)}\right).$
This theorem reduces to a linear problem due to the considerations presented in the Remark 3.1 and the Remark 3.2. By the same covering argument in the proof of Theorem 2.3, it suffices to show the following general theorem for linear equations.
Theorem 3.4 (Hölder Modulus of Continuity - Uniformly Elliptic).
Given $[\lambda,\Lambda]\subseteq(0,\infty)$, there exist $\alpha\in(0,1)$ and $C\geq 1$, such that for $u\in C^{2}(B_{1})$, and $f\in L^{n}(B_{1})$ non-negative with
$\begin{cases}\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u)\leq f\text{ in }B_{1},\\ \mathcal{P}^{+}_{\lambda,\Lambda}(D^{2}u)\geq-f\text{ in }B_{1},\end{cases}$
we have that
$\sup_{r\in(0,1)}r^{-\alpha}\operatorname*{osc}_{B_{r}}u\leq C\left(\operatorname*{osc}_{B_{1}}u+\|f\|_{L^{n}(B_{1})}\right).$
Remark 3.5.
[…]
By the dilation invariance of the Pucci operators, it suffices to show a diminish of oscillation lemma. The details of the argument follow in the same way as it was done for the Laplacian in the proof of Lemma 2.7. The main tool in this case is the homogeneity and invariance by dilations of the Pucci operators.
Lemma 3.6 (Diminish of Oscillation - Uniformly Elliptic).
Given $[\lambda,\Lambda]\subseteq(0,\infty)$, there exist $\delta,\rho,\theta\in(0,1)$, such that for $u\in C^{2}(B_{1})$, $f\in L^{n}(B_{1})$ non-negative with $\|f\|_{L^{n}(B_{1})}\leq\delta$, and
$\begin{cases}\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u)\leq f\text{ in }B_{1},\\ \mathcal{P}^{+}_{\lambda,\Lambda}(D^{2}u)\geq-f\text{ in }B_{1},\end{cases}$
we have that
$\operatorname*{osc}_{B_{1}}u\leq 1\qquad\Rightarrow\qquad\operatorname*{osc}_{B_{\rho}}u\leq(1-\theta).$
3.2.1. The Probabilistic Insight
The original proof of Theorem 2.3, due to Krylov and Safonov [KS79], was originally presented in terms of stochastic processes. In this section we will briefly describe this perspective which motivates the measure estimates in the coming sections.
Consider $u\in C^{1}({\mathbb{R}}^{n})$ satisfying the first-order equation $b\cdot Du=0$ for some vector valued function $b$. The curve $X=X(t)\in C^{1}([0,\infty)\to{\mathbb{R}}^{n})$ is characteristic for the linear and homogeneous PDE if and only if $X^{\prime}=b(X)$, in which case we obtain by the chain rule that $u(X(t))$ remains constant along the trajectory.
We can find a similar connection with second-order PDEs once we allow random dynamics. The simplest case appears when $X(t)$ is a Brownian motion and $u$ is a harmonic function. By Itô’s calculus we get that the expected value $\mathbb{E}[u(X(t))]$ is also constant [^8] And more precisely, one can also say that $u(X(t))$ is a martingale.. We can use this property to compute $u(x)$ by considering a Brownian motion starting at $x$ up to the stopping time when it hits the region where the values of $u$ are prescribed.
Consider $\Omega\subseteq{\mathbb{R}}^{n}$ open and bounded, and $g\in C({\mathbb{R}}^{n}\setminus\Omega)$. If we define $u(x)$ as the expected value of $g(X(\tau))$ where $X(t)$ is a Brownian motion starting at $x$, and $\tau$ is the exit time from $\Omega$, then $u$ satisfies $\Delta u=0$ in $\Omega$ with the complementary condition $u=g$ in ${\mathbb{R}}^{n}\setminus\Omega$ [^9] The continuity of $u$ across the boundary $\partial\Omega$, and the uniqueness of this characterization, depend on the regularity of the boundary $\partial\Omega$; Lipschitz would be enough..
Figure 10. The Harnack inequality provides a lower bound for the probability that a Brownian motion starting at $x\in B_{1/3}$, hits $A\cap B_{\rho}$ before exiting $B_{1}$. This bound depends linearly on the measure of $A\cap B_{\rho}$, but not on the starting point or other geometric characteristics of the set $A$. The figure above illustrates two realizations of the Brownian motion.
Consider $\rho\in(0,1/3)$, a closed set $A$, and let $v_{\rho,A}(x)$ be the probability that a Brownian motion starting at $x$ reaches the set $A\cap\overline{B_{\rho}}$ before exiting $B_{1}$. Therefore
$\Delta v_{\rho,A}=0\text{ in }B_{1}\setminus(A\cap\overline{B_{\rho}})$
with complementary values
$\begin{cases}v=1\text{ on }A\cap\overline{B_{\rho}},\\ v=0\text{ on }{\mathbb{R}}^{n}\setminus B_{1}.\end{cases}$
The probabilistic Harnack inequality (Figure 10) states that for some constant $c>0$ depending only on the dimension, it holds that
$\min_{\overline{B_{1/3}}}v_{\rho,A}\geq c|A\cap B_{\rho}|.$
In particular, if $u$ is a non-negative super-harmonic function and we let $A=\{u\geq 1\}$, we get by the comparison principle that $u\geq v_{\rho,A}$ and then
$\min_{\overline{B_{1/3}}}u\geq c|\{u\geq 1\}\cap B_{\rho}|.$
Notice that the same inequality can also be deduced from Lemma 2.5 and the Markov inequality. Moreover, the proof of the diminish of oscillation (Lemma 2.6) can also be achieved if we consider instead the alternatives whether |$|\{u\geq 1/2\}\cap B_{\rho}|$ is at least half the measure of $B_{\rho}$, or is less than half the measure of the same ball.
This probabilistic approach to the Harnack principle can actually be extended to general diffusions. Given that the identities now relate the distribution of $u$ with point-wise values, these inequalities are known as measure estimates. They replace the mean value property for uniformly elliptic equations without divergence structure.
3.2.2. The Geometric Insight
Another significant insight to the theory was provided by the Minkowski problem. It consist on constructing a convex body with prescribed Gaussian curvature over its boundary. In the case that the boundary is parameterized by the graph of the convex function $u$ we encounter the PDE
$\det(D^{2}u)=K(x)(1+|Du|^{2})^{(n+2)/2}.$
In this context, the leading term is referred to as the Monge-Ampère operator and holds important connections with the field of optimal transport. Notice that it remains elliptic under the assumption of convexity, however it lacks uniform ellipticity.
A key geometrical observation for these problems is the Aleksandrov’s lemma [Ale60, Ale65]. Similar to the mean value property, it allows to control the point-wise values of $u$ in terms of the integral of a second order operator. Notice however that in this scenario the natural integration theorem is no longer the divergence theorem, but the area formula or the formula of change of variables.
Lemma 3.7.
There exists a constant $C\geq 1$ depending only on the dimension such that the following holds: Let $\Omega\subseteq{\mathbb{R}}^{n}$ be an open and convex set and $u\in C^{2}(\Omega)\cap C(\overline{\Omega})$ a convex function with $u=0$ on $\partial\Omega$, then
$\sup_{x\in\Omega}\frac{|u(x)|^{n}}{\operatorname*{dist}(x,\partial\Omega)}\leq C\operatorname*{diam}(\Omega)^{n-1}\int_{\Omega}\det(D^{2}u)$
We suggest that the reader attempt the proof of this lemma after reading this and the next section. See Problem 9 and its hints at the end.
The techniques presented next could be considered refined versions of the convex geometry arguments developed by Aleksandrov, Bakelmann, and Pucci to establish a maximum principle for linear equations without divergence structure. We recommend the lecture [Naz05] for a complete survey on the subject.
3.2.3. Touching Solutions With a Family of Test Functions
One of the main tools available for non-divergence elliptic equations is the comparison principle. In the case of a super-solution $u$, a test function touching $u$ from below inherits information from the equation at the contact point. If we now consider a whole family of test functions, it is reasonable to expect that the corresponding contact set must also be constrained by the equation.
To fix ideas, consider a family of test functions obtained by translations of a given profile
$\varphi_{y_{0}}(x):=\varphi_{0}(x-y_{0}).$
We will call $y_{0}$ the center of the function $\varphi_{y_{0}}$.
For a given function $u\in C(\Omega)$, consider the contact set as $y_{0}$ varies in some set $B$ as
$A:=\bigcup_{y_{0}\in B}\operatorname*{argmin}(u-\varphi_{y_{0}}).$
(Figure 11).
Figure 11. The contact set for a function $u$ is the set of points in the domain that admit a lower supporting graph of the form $\varphi_{y_{0}}+c$.
On one hand, the set $A$ captures important information about $u$. For instance, if $0\in\Omega$, $\varphi_{0}\leq 0$, and $B=B_{r}$, then we get that in the contact set $A$ the values of $u$ are bounded in terms of the radius $r$ and $u(0)$. Indeed, for any $x_{0}\in\operatorname*{argmin}(u-\varphi_{y_{0}})$
[…]
On the other hand, the uniform ellipticity can be used to give a lower bound on the measure of the contact set. For this we notice first that if the paraboloid $\varphi_{y_{0}}$ touches $u$ from below at $x_{0}$, then
$Du(x_{0})=D\varphi_{y_{0}}(x_{0})=D\varphi_{0}(x_{0}-y_{0}).$
If $D\varphi_{0}$ has an inverse, then one can compute the center of the test function touching $u$ from below at $x_{0}$ as $y_{0}=T(x_{0}):=x_{0}-D\varphi_{0}^{-1}(Du(x_{0}))$.
If $T\colon A\to B$ is surjective and Lipschitz, we get by the area formula (change of variables) that
$|B|\leq\int_{A}|\det(DT)|.$
Notice that $DT$ ultimately depends on the Hessian of $u$. Using the previous identity, we observe that the measure of the contact set $A$ can be compared with the measure of $B$, provided some bound on the Hessian of $u$ over $A$.
We can control the eigenvalues of the Hessian of $u$ over $A$ using the contact with the test function and uniform ellipticity: The negative eigenvalues of $D^{2}u$ are automatically bounded by the eigenvalues of $D^{2}\varphi_{y_{0}}$ by the second derivative test. Then we can use that $\mathcal{P}_{\lambda,\Lambda}^{-}(D^{2}u)=\sum_{e\in\operatorname{eig}(D^{2}u)}\lambda e_{+}-\Lambda e_{-}$, to bound the positive eigenvalues of $D^{2}u(x_{0})$ in terms of $\mathcal{P}_{\lambda,\Lambda}^{-}(D^{2}u)$ and the already controlled negative eigenvalues.
3.2.4. Measure Estimates (ABP-type Lemmas)
The following lemma provides a concrete application of this idea when $\varphi_{0}(x)=-\frac{1}{2}|x|^{2}$ is a concave paraboloid, and therefore $T(x)=x+Du(x)$. This approach is know as the method of the sliding paraboloids and was originally developed in [Cab97] and [Sav07].
Lemma 3.8 (Measure Estimate).
Given $[\lambda,\Lambda]\subseteq(0,\infty)$, there exist $\delta,\theta,\eta\in(0,1)$, such that for $u\in C^{2}(B_{1})$ non-negative with $\|(\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u))_{+}\|_{L^{n}(B_{1})}\leq\delta$, we have that
$u(0)\leq\theta\qquad\Rightarrow\qquad|\{u\leq 1\}\cap B_{1}|\geq\eta.$
Remark 3.9.
By homogeneity and dilation invariance we get that under the same hypotheses
$x_{0}\in\{u\leq\theta/2\}\cap\overline{B_{1/2}}\qquad\Rightarrow\qquad|\{u\leq 1/2\}\cap B_{1}|\geq|\{u\leq 1/2\}\cap B_{1/2}(x_{0})|\geq 2^{-n}\eta.$
The contrapositive of this observation gets closer to the key step in the diminish of the oscillation presented in Lemma 3.6. Unfortunately, it is not sufficient to finish the proof.
Assume that $u$ takes values in $[0,1]$, and we want to consider in which direction the oscillation improves, depending on which one of the following measures is larger, |$|\{u\geq 1/2\}\cap B_{1}|$ or |$|\{u\leq 1/2\}\cap B_{1}|$. This would require to have $2^{-n}\eta$ to be at least half the measure of $|B_{1}|$, which is not the case.
Proof of Lemma 3.8.
Consider the family of paraboloids
$\varphi_{y_{0}}(x):=-\frac{1}{2}|x-y_{0}|^{2}+\frac{1}{2}\left(\frac{3}{4}\right)^{2}\leq 1.$
Letting $\theta:=1/4$, we get that for every $y_{0}\in B_{1/4}$ the following holds
$\begin{cases}\varphi_{y_{0}}(0)>\theta,\\ \varphi_{y_{0}}<0\text{ in }{\mathbb{R}}^{n}\setminus B_{1}.\end{cases}$
These properties and the hypotheses $\inf_{B_{1}}u\geq 0$ and $u(0)\leq\theta$, imply that for every $y_{0}\in B_{1/4}$
(3.5) $\emptyset\neq\operatorname*{argmin}(u-\varphi_{y_{0}})\subseteq\{u\leq 1\}\cap B_{1}.$
(Figure 12). Therefore, the result would follow from a lower bound on the measure of the contact set $A=\bigcup_{y_{0}\in B_{1/4}}\operatorname*{argmin}(u-\varphi_{y_{0}})$.
Figure 12. The paraboloids are arranged such that for $y_{0}\in B_{1/4}$ it happens that $\varphi_{y_{0}}(0)>\theta$, and $\varphi_{y_{0}}<0$ in ${\mathbb{R}}^{n}\setminus B_{1}$. Hence a positive function that dips below $\theta$ at the origin must necessarily cross all of these paraboloids. By translating each paraboloid downwards we obtain a contact point in $\{u\leq 1\}\cap B_{1}$.
Consider now the mapping $T\colon A\to B_{1/4}$ defined such that $T(x_{0})=x_{0}+Du(x_{0})$. Equivalently, $y_{0}=T(x_{0})$ is the center of the paraboloid $\varphi_{y_{0}}$ such that $x_{0}\in\operatorname*{argmin}(u-\varphi_{y_{0}})$. By (3.5) it also follows that $T$ is surjective. We finally observe that $DT=I+D^{2}u\geq 0$ over the contact set $A$, just by definition of the contact set and the second derivative test.
Now we will use the uniform ellipticity over the contact set. For every $x_{0}\in A$ we already have that the eigenvalues of $D^{2}u(x_{0})$ are bounded from below by $-1$, then for any positive eigenvalue $e\in\operatorname{eig}(D^{2}u(x_{0}))$ we must have that
$\lambda e\leq(n-1)\Lambda+\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u(x_{0})).$
This means that $\det(D^{2}u+I)\leq C(1+(\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u(x_{0})))_{+}^{n})$ in $A$ for some $C\geq 1$ depending on the dimension $n$ and the ellipticity constants.
By the area formula and the hypothesis $\|(\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u))_{+}\|_{L^{n}(B_{1})}\leq\delta$,
$|B_{1/4}|\leq\int_{A}\det(D^{2}u+I)\leq C|A|+C\delta^{n}.$
By choosing $\delta$ sufficiently small we can absorb the term $C\delta^{n}$ on the left-hand side and get the desired estimate for $\eta:=|B_{1/4}|/(2C)$. ∎
The following lemma uses once again whole family of test functions from below and their corresponding contact set. Its purpose is to localize the measure estimate from the previous result.
Lemma 3.10 (Localization).
Given $[\lambda,\Lambda]\subseteq(0,\infty)$ and $\rho\in(0,1/2)$, there exist $\delta\in(0,1)$ and $M\geq 1$, such that for $u\in C^{2}(B_{1})$ non-negative with $\|(\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u))_{+}\|_{L^{n}(B_{1})}\leq\delta$, we have that
$\min_{\overline{B_{1/2}}}u\leq 1\qquad\Rightarrow\qquad\min_{\overline{B_{\rho}}}u\leq M.$
Figure 13. For any $B_{r}(x_{0})$ in the domain of the equation, we have that if the set $\{u\leq 1\}$ enters $\overline{B_{r/2}(x_{0})}$, then the set $\{u\leq M\}$ has positive density in $B_{\rho r}(x_{0})$. In particular, this prevents the intermediate set $\{1<u\leq M\}$ being too thin, which favours the continuity of $u$.
By combining the measure estimate (Lemma 3.8), the localization (Lemma 3.10), and using the symmetries of the equation, we obtain the following corollary. In Figure 13 we illustrate a rescaled version of this result that will be used in the next section.
Corollary 3.11 (Localized Measure Estimate).
Given $[\lambda,\Lambda]\subseteq(0,\infty)$ and $\rho\in(0,1/2)$, there exist $\delta,\eta\in(0,1)$ and $M\geq 1$, such that for $u\in C^{2}(B_{1})$ non-negative such that $\|(\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u))_{+}\|_{L^{n}(B_{1})}\leq\delta$, we have that
$\min_{\overline{B_{1/2}}}u\leq 1\qquad\Rightarrow\qquad\frac{|\{u\leq M\}\cap B_{\rho}|}{|B_{\rho}|}\geq\eta.$
Proof of Lemma 3.10.
Assume by contradiction that $\min_{\overline{B_{1/2}}}u\leq 1$ but $\min_{\overline{B_{\rho}}}u>M$ for some $M\geq 1$ to be conveniently fixed.
Consider the family of test functions (Figure 14)
$\varphi_{y_{0}}(x):=C_{0}\frac{q(|x-y_{0}|)-q(1-\rho/2)}{q(1/2+\rho/2)-q(1-\rho/2)},\qquad q(r):=\min(r^{-\alpha},(\rho/2)^{-\alpha}),$
where $C_{0},\alpha\geq 1$ will be chosen sufficiently large.
The construction of these functions guarantee that for any $y_{0}\in B_{\rho/2}$, we have that $\varphi_{y_{0}}\geq 1$ in $\overline{B_{1/2}}$, and $\varphi_{y_{0}}<0$ in $\partial B_{1}$. Let us declare now $M:=\sup_{{\mathbb{R}}^{n}}\varphi_{y_{0}}$.
By these properties and the hypotheses $\inf_{B_{1}}u\geq 0$, $\min_{\overline{B_{\rho}}}u>M$, and $\min_{\overline{B_{1/2}}}u\leq 1$ we obtain that for every $y_{0}\in B_{\rho/2}$
(3.6) $\emptyset\neq\operatorname*{argmin}(u-\varphi_{y_{0}})\subseteq B_{1}\setminus\overline{B_{\rho}}.$
Figure 14. The test functions are arranged such that for $y_{0}\in B_{\rho/2}$, they are smooth in $B_{1}\setminus\overline{B_{\rho}}$, remain above one in $\overline{B_{1/2}}$, and below zero in ${\mathbb{R}}^{n}\setminus B_{1}$. Hence, a positive function that goes above their maximum $M$ in $\overline{B_{\rho}}$, and dips below one in $\overline{B_{1/2}}$ must necessarily cross all the functions in $B_{1}\setminus\overline{B_{\rho}}$. As in Lemma 3.8, we can translate the test functions downwards until the touch $u$ at some point in $B_{1}\setminus\overline{B_{\rho}}$.
Let $A=\bigcup_{y_{0}\in B_{\rho/2}}\operatorname*{argmin}(u-\varphi_{y_{0}})$, and $T:A\to B_{\rho/2}$ such that $T(x_{0})=y_{0}$ if $x_{0}\in\operatorname*{argmin}(u-\varphi_{y_{0}})$. This mapping is surjective due to (3.6) and gets computed from the tangency relation at the contact point
$T(x_{0})=x_{0}-D\varphi_{0}^{-1}(Du(x_{0})).$
Here we are using that $D\varphi_{0}:{\mathbb{R}}^{n}\setminus\overline{B_{\rho/2}}\to{\mathbb{R}}^{n}$ is injective, which can be verified by inspection (we will also compute the Hessian of $\varphi_{0}$ in a moment).
By the area formula
(3.7) $|B_{\rho/2}|\leq\int_{A}|\det(DT)|=\int_{A}|\det(I-\Phi^{-1}D^{2}u)|,$
where $\Phi(x):=D^{2}\varphi_{0}(x-T(x))\in{\mathbb{R}}^{n\times n}_{\text{sym}}$.
[…]
$|\det(I-N_{1}^{-1}N_{2})|$ $=\sqrt{\det((I-N_{1}^{-1}N_{2})^{T}(I-N_{1}^{-1}N_{2}))}$
$\leq|I-N_{2}N_{1}^{-1}-N_{1}^{-1}N_{2}+N_{2}N_{1}^{-2}N_{2}|_{\text{op}}^{n/2}$
$\leq(1+2|N_{1}^{-1}|_{\text{op}}|N_{2}|_{\text{op}}+|N_{1}^{-1}|_{\text{op}}^{2}|N_{2}|_{\text{op}}^{2})^{n/2}$
$\leq C(1+|N_{1}^{-1}|_{\text{op}}^{n}|N_{2}|_{\text{op}}^{n})$
$=C\left(1+\left(\frac{\max_{e\in\operatorname{eig}(N_{2})}|e|}{\min_{e\in\operatorname{eig}(N_{1})}|e|}\right)^{n}\right).$
We will need in this way an upper bound on the absolute value of the eigenvalues of $N_{2}=D^{2}u$, and a bound that separates the eigenvalues of $N_{1}=\Phi$ from zero.
For $z\in{\mathbb{R}}^{n}\setminus\overline{B_{\rho/2}}$ we can compute explicitly
$D^{2}\varphi_{0}(z)=C_{0}C_{1}\alpha|z|^{-\alpha-2}\left((\alpha+2)\frac{z\otimes z}{|z|^{2}}-I\right),\qquad C_{1}:=(q(1/2+\rho/2)-q(1-\rho/2))^{-1}.$
From this we get that the eigenvalues of $D^{2}\varphi_{0}(z)$ are: $C_{0}C_{1}\alpha|z|^{-\alpha-2}(\alpha+1)$ with single multiplicity, and $-C_{0}C_{1}\alpha|z|^{-\alpha-2}$ with multiplicity $(n-1)$.
At this point we notice that at any $x_{0}\in A$ and $y_{0}=T(x_{0})$
$\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u(x_{0}))\geq\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}\varphi_{y_{0}}(x_{0}))=C_{0}C_{1}\alpha|x_{0}-y_{0}|^{-\alpha-2}(\lambda(\alpha+1)-\Lambda(n-1)).$
We can make the right-hand side positive by choosing $\alpha:=\Lambda n/\lambda$. Then we can also make the right-hand side $\geq 1$ by fixing $C_{0}$ large enough, depending on $\rho$ and $\alpha$. In conclusion, we have fixed $\alpha$ and $M$ such that we enforce $\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u(x_{0}))\geq 1$ over the contact set $A$. We will see in a moment that this leads to a contradiction.
Once we have fixed the values of $\alpha$ and $M$ we notice that $|D^{2}\varphi_{y_{0}}|_{\text{op}}$ and $|(D^{2}\varphi_{y_{0}})^{-1}|_{\text{op}}$ are bounded over $B_{1}\setminus\overline{B_{\rho}}\supseteq A$, uniformly with respect to $y_{0}\in B_{\rho/2}$. The eigenvalues of $D^{2}u(x_{0})$ are bounded from below by $-|D^{2}\varphi_{y_{0}}(x_{0})|_{\text{op}}$. On the other hand, the equation provides an upper bound for the positive eigenvalues $e\in\operatorname{eig}(D^{2}u(x_{0}))$
$\lambda e\leq\Lambda(n-1)|D^{2}\varphi_{y_{0}}(x_{0})|_{\text{op}}+\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u(x_{0})).$
Putting together the information about the eigenvalues of $N_{1}=\Phi=D^{2}\varphi_{y_{0}}$ and $N_{2}=D^{2}u$, and using that $\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u(x_{0}))\geq 1$ in $A$, we conclude that
$|DT|=|\det(I-\Phi^{-1}D^{2}u)|\leq C(\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u(x_{0})))^{n}\text{ in }A.$
Therefore from (3.7) we obtain $|B_{\rho/2}|\leq C\|(\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u))_{+}\|_{L^{n}(B_{1})}^{n}\leq C\delta^{n}$, however this is a contradiction if we finally take $\delta$ sufficiently small. ∎
{exercise}
[Sun Rising Lemma. Given $u\in C((0,1))$ and $m\in{\mathbb{R}}$, consider the contact set with rays of slope $m$ coming from the right (the sunny region)
$A^{m}:=\left\{x\in(0,1)\,:\,u(x)-mx\geq\sup_{y\in(x,1)}(u(y)-my)\right\}.$
We also consider $S^{m}:=(0,1)\setminus A^{m}$ (the shaded region), see Figure 15.]
Show that for $m>0$
$|S^{m}|\leq m^{-1}\operatorname*{osc}_{(0,1)}u.$
Show that for $m>0$,
$\left\{x\in(0,1)\,:\,\limsup_{h\to 0^{+}}\frac{u(x+h)-u(x)}{h}>m\right\}\subseteq S^{m}.$
Actually, the conclusion
|$|\{u^{\prime}>m\}|\leq m^{-1}\operatorname*{osc}_{(0,1)}u,$
has a trivial proof if we assume $u$ to be differentiable.
Figure 15. The lemma states that as the sun rises from the right or east ($m\to\infty$), the length of the shaded region goes to zero in a quantified way ($|S^{m}|\lesssim m^{-1}$).
{exercise}
[Maximum Principles. The techniques used in the previous two lemmas were originally used to prove the Aleksandrov-Bakelman-Pucci Maximum Principle, for this reason they are also known as ABP-type lemmas. In this case the test functions are linear functions parameterized by its gradient.]
Given $[\lambda,\Lambda]\subseteq(0,\infty)$ there exists $C>1$ such that the following holds: Let $u\in C^{2}(B_{1})\cap C(\overline{B_{1}})$ with $\min_{\partial B_{1}}u\geq 0$, and consider the contact set
$A:=\bigcup_{p\in{\mathbb{R}}^{n}}\operatorname*{argmin}(u-\varphi_{p}),\qquad\varphi_{p}(x):=p\cdot x.$
Then
$\max_{\overline{B_{1}}}u_{-}\leq C\|(\mathcal{P}_{\lambda,\Lambda}^{-}(D^{2}u))_{+}\|_{L^{n}(A)}.$
There exists some constant $C\geq 1$ depending on the dimension such that the following holds. Let $\Omega\subseteq{\mathbb{R}}^{n}$ be an open, convex and bounded set, and $u\in C^{2}(\Omega)\cap C(\overline{\Omega})$ a convex function with $u=0$ on $\partial\Omega$. Then, for every $x_{0}\in\Omega$
$|u(x_{0})|^{n}\leq C\operatorname*{diam}(\Omega)^{n-1}\operatorname*{dist}(x_{0},\partial\Omega)\int_{\Omega}\det(D^{2}u),$
This gives a $1/n$-Hölder modulus of continuity for convex functions at the boundary.
Figure 16. Geometrically, $(y_{0},u_{\varepsilon}(y_{0}))$ gives the vertex of a concave paraboloid that touches $u$ from below. Similar to the approximations of the identity, we obtain that $u_{\varepsilon}$ approximates $u$ as $\varepsilon\to 0$, and gains some regularity. This construction is essential in the theory of viscosity solutions.
{exercise}
[Inf/Sup-Convolutions. Given $\Omega\subseteq{\mathbb{R}}^{n}$ open, $u\in C(\Omega)$, and $\varepsilon>0$, we define the inf-convolution of $u$ as $u_{\varepsilon}\colon{\mathbb{R}}^{n}\to{\mathbb{R}}$ such that
$u_{\varepsilon}(y_{0}):=\inf_{\Omega}\left(u-\varphi^{1/\varepsilon}_{y_{0}}\right),\qquad\varphi^{1/\varepsilon}_{y_{0}}(x):=-\frac{1}{2\varepsilon}|x-y_{0}|^{2}.$
See Figure 16.]
Show that if $0<\varepsilon_{1}<\varepsilon_{2}$, then $u_{\varepsilon_{2}}\leq u_{\varepsilon_{1}}\leq u$.
Show that $u_{\varepsilon}$ converges to $u$ locally uniformly.
Show that $u_{\varepsilon}$ is an $\varepsilon^{-1}$-concave function [^11] This means that $x\mapsto u_{\varepsilon}(x)+|x|^{2}/(2\varepsilon)$ is concave. Equivalently, every point $(y_{0},u_{\varepsilon}(y_{0}))$ in the graph of $u_{\varepsilon}$ has a supporting paraboloid of the from $\varphi_{x_{0}}^{-1/\varepsilon}+C$ from above. Notice that in case that the Hessian exists, this notion means that $D^{2}u_{\varepsilon}\leq\varepsilon^{-1}I$. We can also say that a function $u$ is $\alpha$-convex if and only if $-u$ is $(-\alpha)$-concave. In general, these types of functions are known as semi-concave/convex..
Show the semi-group property: $(u_{1/\alpha_{1}})_{1/\alpha_{2}}=u_{1/(\alpha_{1}+\alpha_{2})}$.
One can easily check that
$\{u=u_{\varepsilon}\}\subseteq\bigcup_{y_{0}\in{\mathbb{R}}^{n}}\operatorname*{argmin}(u-\varphi_{y_{0}}^{1/\varepsilon}).$
Unfortunately, the inclusion could be strict in some cases. To get an equality we need an additional construction.
Consider the sup-convolution
$u^{\varepsilon}(y_{0}):=\sup_{\Omega}\left(u-\varphi^{-1/\varepsilon}_{y_{0}}\right).$
Show that $\Gamma_{\varepsilon}:=(u_{\varepsilon})^{\varepsilon}$ is the upper-envelope for the family of paraboloids of the form $\varphi^{1/\varepsilon}_{y_{0}}+C$ that remain below $u$. In particular
$\{u=\Gamma_{\varepsilon}\}=\bigcup_{y_{0}\in{\mathbb{R}}^{n}}\operatorname*{argmin}(u-\varphi_{y_{0}}^{1/\varepsilon}).$
{exercise}
(Measure Estimate with Small Drift) Show that for $[\lambda,\Lambda]\subseteq(0,\infty)$, there exist $\delta,\theta,\eta\in(0,1)$, such that the following holds: Let $u\in C^{2}(B_{1})$, and $b,f\in L^{n}(B_{1})$ with $\|b\|_{L^{n}(B_{1})}\leq\delta$, $\|f\|_{L^{n}(B_{1})}\leq\delta$, all of them non-negative functions such that
$\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u)-b|Du|\leq f\text{ in }B_{1}.$
Then
$u(0)\leq\theta\qquad\Rightarrow\qquad|\{u\leq 1\}\cap B_{1}|\geq\eta.$
(Localization with Small Drift) Prove a similar extension for Lemma 3.10 and its corollary.
(ABP Maximum Principle) Prove a similar extension for the ABP Maximum Principle without the smallness condition on $\|b\|_{L^{n}(B_{1})}$. Given $[\lambda,\Lambda]\subseteq(0,\infty)$ there exists $C>1$ such that the following holds: Let $u\in C^{2}(B_{1})\cap C(\overline{B_{1}})$, and $b,f\in L^{n}(B_{1})$ non-negative with $\|b\|_{L^{n}(B_{1})}\leq\Lambda$ such that
$\begin{cases}\mathcal{P}_{\lambda,\Lambda}^{-}(D^{2}u)-b|Du|\leq f\text{ in }B_{1},\\ u\geq 0\text{ on }\partial B_{1}.\end{cases}$
Consider the contact set
$A:=\bigcup_{p\in{\mathbb{R}}^{n}}\operatorname*{argmin}(u-\varphi_{p}),\qquad\varphi_{p}(x):=p\cdot x.$
Then
$\max_{\overline{B_{1}}}u_{-}\leq C\|f\|_{L^{n}(A)}.$
(Localized Measure Estimate) Prove a similar extension for Corollary 3.11 without the smallness condition on $\|b\|_{L^{n}(B_{1})}$. Given $[\lambda,\Lambda]\subseteq(0,\infty)$ and $\rho\in(0,1/2)$, there exist $\delta,\eta\in(0,1)$ and $M\geq 1$, such that the following holds: Let $u\in C^{2}(B_{1})$, and $b,f\in L^{n}(B_{1})$ with $\|b\|_{L^{n}(B_{1})}\leq\Lambda$ and $\|f\|_{L^{n}(B_{1})}\leq\delta$, all of them non-negative functions such that
$\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u)-b|Du|\leq f\text{ in }B_{1}.$
Then
$\min_{\overline{B_{1/2}}}u\leq 1\qquad\Rightarrow\qquad\frac{|\{u\leq M\}\cap B_{\rho}|}{|B_{\rho}|}\geq\eta.$
{exercise}
[Measure Estimates for Parabolic Equations. For dynamic problems, we consider contact points $(x_{0},t_{0})\in B_{1}\times(-1,0]$ with respect to a function $u=u(x,t)\in C(B_{1}\times(-1,0])$ and a test function $\varphi=\varphi(x,t)$ from below. The time $t_{0}$ is the first time when the graph of $\varphi$ touches the graph of $u$ from below, and the point $x_{0}$ is just any point where the contact occurs at time $t_{0}$.]
[…]
$\begin{cases}\partial_{t}u-\mathcal{P}_{\lambda,\Lambda}^{-}(D^{2}u)\geq-f\text{ in }B_{1}\times(-1,0],\\ u\geq 0\text{ on }\partial_{\text{par}}(B_{1}\times(-1,0]).\end{cases}$
Consider for $p\in{\mathbb{R}}^{n}$, $h\in{\mathbb{R}}$, and $\varphi_{p,h}(x):=p\cdot x-h$, the contact set
$A:=\{(x_{0},t_{0})\in B_{1}\times(-1,0]\,:\,\text{For some $(p_{0},h_{0})\in{\mathbb{R}}^{n+1}$,}$
$u>\varphi_{p_{0},h_{0}}$ in $B_{1}\times(-1,t_{0})$,
$\text{and $u(x_{0},t_{0})=\varphi_{p_{0},h_{0}}(x_{0},t_{0})$}\}.$
Then
$\max_{\overline{B_{1}}\times[-1,0]}u_{-}\leq C\|f\|_{L^{n+1}(A)}.$
(Measure Estimate) Given $[\lambda,\Lambda]\subseteq(0,\infty)$ there exist $\delta,\theta,\eta\in(0,1)$ such that the following holds: Let $u\in C^{2}(B_{1}\times(-1,0])$ and $f\in L^{n+1}(B_{1}\times(-1,0])$ with $\|f\|_{L^{n+1}(B_{1}\times(-1,0])}\leq\delta$, both non-negative functions such that
$\partial_{t}u-\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u)\geq-f\text{ in }B_{1}\times(-1,0].$
Then
$u(0,0)\leq\theta\qquad\Rightarrow\qquad|\{u\leq 1\}\cap B_{1}\times(-1,0]|\geq\eta.$
(Localization) Given $[\lambda,\Lambda]\subseteq(0,\infty)$ and $\rho\in(0,1/2)$ there exist $\delta,\eta\in(0,1)$ and $M\geq 1$, such that the following holds: Let $u\in C^{2}(B_{1}\times(-1,0])$ and $f\in L^{n+1}(B_{1}\times(-1,0])$ with $\|f\|_{L^{n+1}(B_{1}\times(-1,0])}\leq\delta$, both non-negative functions such that
$\partial_{t}u-\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u)\geq-f\text{ in }B_{1}\times(-1,0].$
Then
$\min_{\overline{B_{1/2}}\times[-1/2,0]}u\leq 1\qquad\Rightarrow\qquad\min_{\overline{B_{\rho}}\times[-1,-1+\rho]}u\leq M.$
(Localized Measure Estimate) Given $[\lambda,\Lambda]\subseteq(0,\infty)$ and $\rho\in(0,1/2)$ there exist $\delta,\eta\in(0,1)$ and $M\geq 1$, such that the following holds: Let $u\in C^{2}(B_{1}\times(-1,0])$ and $f\in L^{n+1}(B_{1}\times(-1,0])$ with $\|f\|_{L^{n+1}(B_{1}\times(-1,0])}\leq\delta$, both non-negative functions such that
$\partial_{t}u-\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u)\geq-f\text{ in }B_{1}\times(-1,0].$
Then
$\min_{\overline{B_{1/2}}\times[-1/2,0]}u\leq 1\qquad\Rightarrow\qquad\frac{|\{u\leq M\}\cap B_{\rho}\times(-1/2-\rho,-1/2]|}{|B_{\rho}\times(-1/2-\rho,-1/2]|}\geq\eta.$
{exercise}
[Measure Estimates for the Hessian I. We have seen in this section and the previous problems that the contact set captures information about $u$ and also its gradient, due to the first derivative test. It is natural to expect that it also captures information about its Hessian, due to the second derivative test. For $M>0$, $B\subseteq{\mathbb{R}}^{n}$, and $u\in C^{2}(B_{1})$, consider the contact sets
$A^{M}_{B}:=\bigcup_{y_{0}\in B}\operatorname*{argmin}\left(u-\varphi_{y_{0}}^{M}\right)\subseteq\{|(D^{2}u)_{-}|_{\text{op}}\leq M\},\qquad\varphi_{y_{0}}^{M}(x):=-\frac{M}{2}|x-y_{0}|^{2}.$
Given $[\lambda,\Lambda]\subseteq(0,\infty)$ and $\rho\in(0,1/2)$, there exist $\delta,\eta\in(0,1)$ and $M\geq 1$, such that for $u\in C^{2}(B_{1})$ with $\|(\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u))_{+}\|_{L^{n}(B_{1})}\leq\delta$,
$A^{1}_{\{0\}}\cap\overline{B_{1/2}}\neq\emptyset\qquad\Rightarrow\qquad\frac{|A^{M}_{{\mathbb{R}}^{n}}\cap B_{\rho}|}{|B_{\rho}|}\geq\eta.$
3.2.5. Diminish of Distribution
The localized measure estimate proved in the previous section concludes that
|$|\{u>M\}\cap B_{\rho}|\leq(1-\eta)|B_{\rho}|.$
In this section we will iterate this estimate and obtain the following geometric decay for the distribution
|$|\{u>M^{k}\}\cap B_{\rho}|\leq(1-\eta)|\{u>M^{k-1}\}\cap B_{\rho}|\leq\ldots\leq(1-\eta)^{k}|B_{\rho}|.$
As soon as $(1-\eta)^{k}$ becomes less than $1/2$ we are in shape to prove the diminish of oscillation (Lemma 3.6).
By the homogeneity of the equation, it suffices to show the step from $k=0$ to $k=1$
|$|\{u>M\}\cap B_{\rho}|\leq(1-\eta)|\{u>1\}\cap B_{\rho}|,$
which by rearranging the terms is equivalent to the following estimate (Figure 17)
(3.8) |$|\{1<u\leq M\}\cap B_{\rho}|\geq\eta|\{u>1\}\cap B_{\rho}|.$
At this point we pose the equation in $B_{3\sqrt{n}}$, replace $B_{\rho}$ for the cube $Q_{1}=(-1/2,1/2)^{n}$ of unit length, and $\overline{B_{1/2}}$ for $\overline{Q_{3}}=[-3/2,3/2]^{n}$. This modification will be convenient to implement a covering argument in its proof.
Figure 17. The estimate |$|\{1<u\leq M\}\cap B_{\rho_{1}}|\geq\eta|\{u>1\}\cap B_{\rho_{1}}|$ resembles the isoperimetric inequality where the set $\{1<u\leq M\}$ plays the role of the perimeter for the set $\{u>1\}$.
Lemma 3.12 (Diminish of Distribution).
Given $[\lambda,\Lambda]\subseteq(0,\infty)$, there exist $\delta,\eta\in(0,1)$ and $M\geq 1$, such that for $u\in C^{2}(B_{3\sqrt{n}})$ non-negative with $\|(\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u))_{+}\|_{L^{n}(B_{3\sqrt{n}})}\leq\delta$,
$\min_{\overline{Q_{3}}}u\leq 1\qquad\Rightarrow\qquad|\{1<u\leq M\}\cap Q_{1}|\geq\eta|\{u>1\}\cap Q_{1}|.$
By taking $\varepsilon=-\ln(1-\eta)/\ln M$ we obtain the following result with an argument similar to the one in the proof of Lemma 2.7.
Lemma 3.13 (Weak Harnack Inequality - Uniformly Elliptic).
Given $[\lambda,\Lambda]\subseteq(0,\infty)$, there exist $\varepsilon\in(0,1)$ and $C\geq 1$ such that for $u\in C^{2}(B_{3\sqrt{n}})$ non-negative,
$\sup_{\mu>0}\mu^{\varepsilon}|\{u\geq\mu\}\cap Q_{1}|\leq C\left(\min_{\overline{Q_{3}}}u+\|(\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u))_{+}\|_{L^{n}(B_{3\sqrt{n}})}\right)^{\varepsilon}.$
A rough idea to prove the diminish of distribution (Lemma 3.12) would be to add up the information given by the localized measure estimate (Corollary 3.11), rescaled to a collection of disjoint balls used to cover the set $\{u>1\}\cap Q_{1}$. If the ball $B_{r}(x_{0})$, contained in the domain of the equation, is such that $r/2=\operatorname*{dist}(x_{0},\{u\leq 1\})$, we have that $\{u\leq M\}\cap B_{\rho r}(x_{0})=\{1<u\leq M\}\cap B_{\rho r}(x_{0})$ has positive density in $B_{\rho r}(x_{0})$. This means that the estimate (3.8) holds in this local configuration (replacing $B_{\rho}$ by $B_{\rho r}(x_{0})$). To integrate this local pieces of information we can use the dyadic decomposition of an relatively open set of the unit cube.
The dyadic cubes contained in the unit cube $\overline{Q_{1}}=[-1/2,1/2]^{n}$, are a collection of closed cubes constructed by an iterative procedure. Each dyadic cube has the form
$\overline{Q_{2^{-k}}(x_{0})}=\prod_{i=1}^{n}[(x_{0})_{i}-2^{-(k+1)},(x_{0})_{i}+2^{-(k+1)}],$
where $x_{0}\in{\mathbb{R}}^{n}$ is its center, $2^{-k}$ is the length of its sides, and $k$ is a non-negative integer. The unit cube $\overline{Q_{1}}$ is dyadic and forms the zeroth generation. Each successive generation is obtained from the previous one by sub-dividing each cube $Q$ in $2^{n}$ congruent cubes of half the length. We call these cubes the descendants of $Q$, and for each descendant of $Q$ we say that $Q$ is its progenitor.
For a relatively open set $E\subseteq\overline{Q_{1}}$, its dyadic decomposition is the set of dyadic cubes $Q\subseteq E$ such that the progenitor of $Q$ is not fully contained in $E$, see [SS05, Theorem 1.4]. This collection forms an almost-disjoint covering of $E$ (any pair of cubes have disjoints interiors). Figure 18 illustrates a few stages in the decomposition of $\{u>1\}\cap\overline{Q_{1}}$.
Figure 18. Each cube $Q$ in the dyadic decomposition of $E=\{u>1\}\cap\overline{Q_{1}}$ is selected by the following algorithm: Starting from the zeroth generation, a dyadic cube will be discarded (with all its descendants) if it does not intersect $E$, it will be selected if it is contained in $E$, otherwise its descendants will be passed to the next iteration.
Proof of Lemma 3.12.
Let $M>1$ and $\delta_{0},\eta_{0}\in(0,1)$ be the constants from Corollary 3.11 with respect to $\rho=1/(6\sqrt{n})$ and the same ellipticity constants. Assume that the hypotheses of Lemma 3.12 hold for $\delta:=\delta_{0}/(3\sqrt{n})$.
Consider an arbitrary dyadic cube $\overline{Q_{\ell}(x_{0})}$ in the decomposition of $\{u>1\}\cap\overline{Q_{1}}$. Then the ball $\overline{B_{3\ell\sqrt{n}/2}(x_{0})}$ contains the predecessor of $Q$, and as a consequence of the construction, it must intersect the set $\{u\leq 1\}$. By the Corollary 3.11, we get that |$|\{u\leq M\}\cap Q_{\ell}(x_{0})|\geq\eta|Q_{\ell}(x_{0})|$ for some $\eta\in(0,1)$, to be fixed from $\eta_{0}$. From this we will automatically get the desired conclusion, just by summing the pieces,
|$|\{1<u\leq M\}\cap Q_{1}|=\sum_{Q}|\{u\leq M\}\cap Q|\geq\eta\sum_{Q}|Q|=\eta|\{u>1\}\cap Q_{1}|.$
The remain of the proof will justify this claim about $Q_{\ell}(x_{0})$.
For $r=\ell/(2\rho)=3\ell\sqrt{n}$, the rescaling $v(x)=u(rx+x_{0})$ verifies that
$\begin{cases}\|\mathcal{P}_{\lambda,\Lambda}^{-}(D^{2}v)\|_{L^{n}(B_{1})}=r\|\mathcal{P}_{\lambda,\Lambda}^{-}(D^{2}u)\|_{L^{n}(B_{r}(x_{0}))}\leq\delta_{0},\\ \min_{\overline{B_{1/2}}}v=\min_{\overline{B_{3\ell\sqrt{n}/2}(x_{0})}}u\leq 1.\end{cases}$
By the Corollary 3.11, we get that
|$\frac{|\{v\leq M\}\cap B_{\rho}|}{|B_{\rho}|}=\frac{|\{u\leq M\}\cap B_{\ell/2}(x_{0})|}{|B_{\ell/2}(x_{0})|}\geq\eta_{0}.$
Then we can now conclude the desired estimate for the dyadic cube $Q_{\ell}(x_{0})\supseteq B_{\ell/2}(x_{0})$ by setting $\eta:=\eta_{0}|B_{1/2}|$. ∎
Remark 3.14 (Calderón-Zygmund Decomposition).
An alternative argument used in [CC95] consists on selecting the dyadic cubes according to the following criteria in terms of $F=\{u>M\}\cap\overline{Q_{1}}$ and $\eta\in(0,1)$, with $|F|\leq(1-\eta)$. The dyadic cube $Q$ is selected if $|F\cap Q|/|Q|>(1-\eta)$, but none of its predecessors satisfy a similar lower bound for the density of $F$. Due to the Lebesgue differentiation theorem, we get that the selected cubes must cover $F$ up to a set of measure zero.
The contrapositive of Corollary 3.11 then implies that the predecessor of every selected cube must be contained in $E=\{u>1\}\cap\overline{Q_{1}}$, hence if we consider the covering given by the predecessors of the cubes selected by the algorithm we get that
$|F|\leq\sum_{Q}|F\cap Q|\leq(1-\eta)\sum_{Q}|Q|\leq(1-\eta)|E|.$
Figure 19. For $x_{0}\in\{u>1\}$, and $r/2=\operatorname*{dist}(x_{0},\{u\leq 1\})$, we get that the set $\{1<u\leq M\}$ has positive density in $B_{\rho r}(x_{0})$. The idea is then to add up this local information. The caveat is that the ball $B_{\rho r}(x_{0})$ may spill out of $B_{\rho}$.
Growing Ink-spots. It is also possible to prove the diminish of distribution (Lemma 3.12) using Vitali’s covering lemma (Figure 19). The following result provides an alternative proof to the Lemma 3.12.
Lemma 3.15 (Growing Ink-spots).
Let $\rho_{0}=1/6$ and $\rho_{1}=1/7$. Given $\eta\in(0,1)$ and $F\subseteq E\subseteq{\mathbb{R}}^{n}$ such that $\overline{B_{\rho_{1}}}\cap F\neq\emptyset$, and for every ball $B_{r}(x_{0})\subseteq B_{1}$
$\overline{B_{r/2}(x_{0})}\cap F\neq\emptyset\qquad\Rightarrow\qquad\frac{|E\cap B_{\rho_{0}r}(x_{0})|}{|B_{\rho_{0}r}(x_{0})|}\geq\eta,$
then
$|(E\setminus F)\cap B_{\rho_{1}}|\geq 5^{-n}\eta|B_{\rho_{1}}\setminus F|.$
Notice that the hypothesis in the previous lemma is the one satisfied by the sets $F=\{u\leq 1\}$ and $E=\{u\leq M\}$ in Figure 13. This makes Lemma 3.12 a corollary of the previous claim. The following proof appeared in [Sav07].
Proof.
Consider for every $x_{0}\in B_{\rho_{1}}\setminus F$ the radius $r(x_{0}):=2\operatorname*{dist}(x_{0},F)$. Given that $F$ barely meets $B_{r(x_{0})/2}(x_{0})$, we have that $E$ has positive density in $B_{\rho_{0}r(x_{0})}(x_{0})$, provided that $B_{r(x_{0})}(x_{0})\subseteq B_{1}$. The collection $\{B_{\rho_{0}r(x_{0})}(x_{0})\}$ covers $B_{\rho_{1}}\setminus F$. Although these balls may not be disjoint, we can still appeal to Vitali’s Covering Lemma to get a disjoint set of balls that covers a fixed fraction of $B_{\rho_{1}}\setminus F$. A second problem arises because the balls $B_{\rho_{0}r(x_{0})}(x_{0})$ are not necessarily contained in $B_{\rho_{1}}$.
To address this second issue we consider a point $x_{1}\in\overline{B_{\rho_{1}}}\cap\partial B_{r(x_{0})/2}(x_{0})$, on the line going through 00 and $x_{0}$. In the exceptional case $x_{0}=0$, we take $x_{1}=0$. Then, for the midpoint $x_{2}:=(x_{0}+x_{1})/2$, we get that $B_{r(x_{0})/4}(x_{2})\subseteq B_{r(x_{0})/2}(x_{0})\cap B_{\rho_{1}}$ (Figure 20).
The collection $\{\overline{B_{r(x_{0})/4}(x_{2})}\}$ covers $B_{\rho_{1}}\setminus F$. Because $B_{3r(x_{0})/4}(x_{2})\supseteq B_{r(x_{0})/2}(x_{0})$ we also get that $\overline{B_{3r(x_{0})/4}(x_{2})}\cap F\neq\emptyset$. Notice that the ratio
$\frac{r(x_{0})/4}{2\times 3r(x_{0})/4}=\frac{1}{6}=\rho_{0}$
was conveniently chosen to fix this construction. In other words, as $F$ meets $B_{3r(x_{0})/4}(x_{2})$, we have by hypothesis that $|E\cap B_{r(x_{0})/4}(x_{2})|\geq\eta|B_{r(x_{0})/4}|$.
Figure 20. For $x_{0}\in B_{\rho_{1}}\setminus F$ and $r=r(x_{0})$, we may find the problem that $B_{r/2}(x_{0})$ spills outside $B_{\rho_{1}}$. We fix this problem by constructing $B_{r/4}(x_{2})\subseteq B_{r/2}(x_{0})\cap B_{\rho_{1}}$. As $F$ enters $B_{r/2}(x_{0})$, then it must also enter $B_{3r/4}(x_{2})$, and by hypothesis we obtain that $E$ must have positive density in $B_{r/4}(x_{2})$.
Finally, the hypothesis $\overline{B_{\rho_{1}}}\cap F\neq\emptyset$ implies that $r(x_{0})\leq 4\rho_{1}$. Then $B_{3r(x_{0})/2}(x_{2})\subseteq B_{7\rho_{1}}\subseteq B_{1}$ by the convenient choice of the radius $\rho_{1}=1/7$.
Let $\mathcal{B}\subseteq\{B_{r(x_{0})/4}(x_{2})\}$ be a countable collection of disjoint balls such that
$\sum_{B\in\mathcal{B}}|B|\geq 5^{-n}|B_{\rho_{1}}\setminus F|.$
We obtain the desired estimate by using that $|E\cap B|\geq\eta|B|$. ∎
{exercise}
[Dyadic Covering Lemma for Parabolic Equations. The localized measure estimate in Problem 12 has a shift in time in relation to the sets where $u$ dips below one ($\overline{B_{1/2}}\times[-1/2,0]$), and the set where the density of $\{u\leq M\}$ gets bounded from below ($B_{\rho}\times(-1/2-\rho,-1/2]$). This requires a non-trivial modification in the dyadic covering argument to get the corresponding weak Harnack inequality. We define the dyadic cylinders generated from $\overline{Z_{1}}:=\overline{Q_{1}}\times[-1,0]$, with a similar procedure that sub-divides the spatial lengths in halves and the temporal lengths in fourths.]
(Staking Lemma) For every cylinder $Z=Q\times[t-r,t]$ and $m\geq 1$, let $Z^{m}=Q\times[t,t+mr]$. Show that for any collection $\mathcal{Z}$ of dyadic cylinders,
$\left|\bigcup_{Z\in\mathcal{Z}}Z^{m}\right|\geq\frac{m}{m+1}\left|\bigcup_{Z\in\mathcal{Z}}(Z^{m}\cup Z)\right|.$
A complete proof of the weak Harnack inequality for uniformly elliptic parabolic equations requires a few more technical steps. A complete treatment can be consulted in [Wan92] and the notes [IS12]. A shorter argument is also available using a parabolic version of Vitali’s lemma, it can be consulted in [Wan13] and the notes [Moo17].
{exercise}
[Measure Estimates for the Hessian II. The diminish of distribution strategy can also be used on the distribution for the second derivative estimates stated in Problem 13. Due to scaling considerations, we introduce the Hardy–Littlewood maximal operator of a function $f$, denoted as
$m(f)(x):=\sup_{r>0}\fint_{B_{r}(x)}|f|.$
Given $[\lambda,\Lambda]\subseteq(0,\infty)$, there exist $\delta,\eta\in(0,1)$ and $M\geq 1$, such that the following holds: Let $u\in C^{2}(B_{3\sqrt{n}})$ with $\|(\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u))_{+}\|_{L^{n}(B_{3\sqrt{n}})}\leq\delta$ and $A^{1}_{\{0\}}\cap\overline{Q_{3}}\neq\emptyset$. Then, for any cube $Q$ in the dyadic decomposition of the set
$((Q_{1}\setminus A^{1}_{{\mathbb{R}}^{n}})\cup\{m((\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u))_{+}^{n})\geq\delta^{n}\})\cap Q_{1}$
we have that $|Q\cap A^{M}_{{\mathbb{R}}^{n}}|\geq\eta|Q|$.
Given $[\lambda,\Lambda]\subseteq(0,\infty)$, there exist $\varepsilon\in(0,1)$ and $C\geq 1$ such that for $u\in C^{2}(B_{3\sqrt{n}})$
$\sup_{\mu>0}\mu^{\varepsilon}|\{|(D^{2}u)_{-}|_{\text{op}}\geq\mu\}\cap Q_{1}|\leq C\left(\operatorname*{osc}_{B_{3\sqrt{n}}}u+\|(\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u))_{+}\|_{L^{n}(B_{3\sqrt{n}})}\right)^{\varepsilon}.$
3.2.6. Summary of the Proofs
At this point one already have all the results required to prove the interior Hölder estimate for uniformly elliptic equation in non-divergence form, Theorem 3.4. We have skipped some details in order to highlight the fundamental ideas. To gain a better view of the theory let us recapitulate the main steps, and compare them with the analogous results for the Laplacian. Finally we quickly assemble the missing proofs.
Measure Estimates: For the Laplacian it is represented by the mean value property, proved as an application of the divergence theorem. For equations without divergence structure, we proved in Lemma 3.8 and Lemma 3.10 that for a super-solution, the measure of the lower contact set with a family of tests functions is bounded from below. It relies in a different integration theorem, the area formula. These proofs are the only ones where the uniform ellipticity is actually used. Corollary 3.11 combines Lemma 3.8 and Lemma 3.10 to obtain a measure estimate localized at the ball $B_{\rho}$.
Proof of Corollary 3.11.
Let $M_{0}$ be the constant from the localization Lemma 3.10 with respect to the radius $\rho_{0}=\rho/2$. By this lemma we get that $u(x_{0})\leq M_{0}$ for some $x_{0}\in B_{\rho/2}$.
Let $\theta$ and $\eta_{0}$ from the Lemma 3.8. By applying this lemma to $\theta M_{0}^{-1}u((\rho/2)x+x_{0})$, we get that |$|\{u\leq\theta^{-1}M_{0}\}\cap B_{\rho/2}(x_{0})|\geq\eta_{0}(\rho/2)^{n}$. Therefore the desired result follows for $M:=\theta^{-1}M_{0}$ and $\eta:=\eta_{0}/|B_{2}|$. ∎
Weak Harnack Inequality: For the Laplacian it followed from the mean value property by a geometric observation (Figure 3). In the more general case, we found a way to iterate the measure estimate (Corollary 3.11) in order to get a geometric decay of the distribution |$|\{u>M^{k}\}\cap Q_{1}|$ in Lemma 3.12.
Proof of Lemma 3.13.
Follows as in the proof of Lemma 2.7, assuming first that for $\delta\in(0,1)$ from Lemma 3.12
$\begin{cases}\|(\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u))_{+}\|_{L^{n}(B_{3\sqrt{n}})}\leq\delta,\\ \min_{\overline{Q_{3}}}u\leq 1.\end{cases}$
Otherwise, we use the homogeneity of the Pucci operator and apply the following reasoning to $u/\left(\min_{\overline{Q_{3}}}u+\delta^{-1}\|(\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u))_{+}\|_{L^{n}(B_{3\sqrt{n}})}\right)$.
For $M$ and $\eta$ as in Lemma 3.12, it follows inductively and once again by the homogeneity of the operator, that for every integer $k\geq 1$,
|$|\{u>M^{k}\}\cap Q_{1}|\leq(1-\eta)^{k}.$
The inductive step follows by applying Lemma 3.12 to $M^{-k}u$. Notice that for $\varepsilon:=-\ln(1-\eta)/\ln M$, we can replace $(1-\eta)^{k}$ in the right-hand side with $(M^{k})^{-\varepsilon}$.
Given $\mu>1$, let $k$ be a non-negative integer such that $M^{k}\leq\mu<M^{k+1}$. Then, once we fix $C:=M^{\varepsilon}$, we obtain
|$|\{u>\mu\}\cap Q_{1}|\leq|\{u>M^{k}\}\cap Q_{1}|\leq(1-\eta)^{k}=C(M^{k+1})^{-\varepsilon}\leq C\mu^{-\varepsilon}.$
If $\mu\in(0,1]$, the desired estimate on the distribution is trivially true. ∎
Diminish of the Oscillation: The proof of the interior Hölder estimate (Theorem 3.4) follows from the dilation invariance of the equation, in the same way as for the Laplacian (proof of the Lemma 2.7), but using the diminish of oscillation Lemma 3.6 instead. Lemma 3.6 relies on the weak Harnack inequality (Lemma 3.13).
Proof of Lemma 3.6.
Assume without loss of generality that $0\leq u\leq 1$ in $B_{1}$. If |$|\{u\geq 1/2\}\cap Q_{1/(3\sqrt{n})}|\geq|Q_{1/(3\sqrt{n})}|/2$, we apply a rescaling of Lemma 3.13 to get that $u\geq\theta:=C^{-1}2^{-(1+2\varepsilon)/\varepsilon}$ in $B_{1/(6\sqrt{n})}\subseteq Q_{1/(3\sqrt{n})}$, provided that $\delta:=C^{-1}2^{-(1+2\varepsilon)/\varepsilon}$. Otherwise we apply Lemma 3.13 to $(1-u)$ to get a similar improvement of the upper bound of $u$. ∎
{exercise}
[Fractional Laplacian. Elliptic integro-differential operators, such as the fractional Laplacians, present in their own definition some sort of mean value property. The present problem aims to illustrate this claim from the comparison principle. For $\sigma\in(0,2)$ we define the fractional Laplacian $\Delta^{\sigma/2}=-(-\Delta)^{\sigma/2}$ in terms of its Fourier symbol $\widehat{\Delta^{\sigma/2}}=-|\xi|^{\sigma}$. From this definition we can evaluate the operator on $u\in L^{1}({\mathbb{R}}^{n})$ at a point $x$ where $u$ is regular enough ($C^{2}$ is enough)
[…]
The constant $C$ appearing above is a positive number that depends only on the parameter $\sigma$ and the dimension $n$.]
(Comparison Principle) Show that for every $u,v\in L^{1}({\mathbb{R}}^{n})$ such that $u\leq v$ and $u(x_{0})=v(x_{0})$ it holds that $\Delta^{\sigma/2}u(x_{0})\leq\Delta^{\sigma/2}v(x_{0})$.
(Weak Harnack) Show that for every $\sigma\in(0,2)$ there exists $C\geq 1$ such that for $u\in L^{1}({\mathbb{R}}^{n})\cap C^{2}(B_{3/4})$ non-negative with $\|(\Delta^{\sigma/2}u)_{+}\|_{L^{\infty}(B_{3/4})}\leq 1$
$\inf_{B_{1/2}}u\leq 1\qquad\Rightarrow\qquad\int_{B_{1}\setminus B_{3/4}}u\leq C.$
(Hölder Estimate) Show that for every $\sigma\in(0,2)$ there exist $\alpha\in(0,1)$ and $C\geq 1$ such that the following Hölder estimate holds for $u\in L^{1}({\mathbb{R}}^{n})\cap C^{2}(B_{3/4})$
$\sup_{r\in(0,1)}r^{-\alpha}\operatorname*{osc}_{B_{r}}u\leq C\left(\operatorname*{osc}_{B_{1}}u+\|u\|_{L^{1}({\mathbb{R}}^{n}\setminus B_{1})}+\|\Delta^{\sigma/2}u\|_{L^{\infty}(B_{3/4})}\right).$
{exercise}
Show that for $[\lambda,\Lambda]\subseteq(0,\infty)$, there exist $\alpha\in(0,1)$ and $C\geq 1$, such that the following holds: Let $F\in C({\mathbb{R}}^{n\times n}_{\text{sym}})$ be translation invariant and uniformly elliptic with respect to $[\lambda,\Lambda]$, $\Omega\subseteq{\mathbb{R}}^{n}$ open, and $u\in C^{3}(\Omega)$ with $F(D^{2}u)=0$ in $\Omega$. Then
$[Du]^{(1+\alpha)}_{C^{0,\alpha}(\Omega)}\leq C\operatorname*{osc}_{\Omega}u.$
Remark 3.16.
[…]
{exercise}
Show that for $[\lambda,\Lambda]\subseteq(0,\infty)$ and $\varepsilon>0$, there exists some $C\geq 1$ such that the following maximum principle holds for $u\in C^{2}(B_{1})$
$\max_{\overline{B_{1/2}}}u_{+}\leq C\left(\|u_{+}\|_{L^{\varepsilon}(B_{1})}+\|(\mathcal{P}^{+}_{\lambda,\Lambda}u)_{-}\|_{L^{n}(B_{1})}\right).$
{exercise}
Show that for $[\lambda,\Lambda]\subseteq(0,\infty)$ there exists some $C\geq 1$ such that for $u\in C^{2}(B_{1})$, and $f\in L^{n}(B_{1})$, both non-negative functions with
$\begin{cases}\mathcal{P}_{\lambda,\Lambda}^{-}(D^{2}u)\leq f\text{ in }B_{1},\\ \mathcal{P}_{\lambda,\Lambda}^{+}(D^{2}u)\geq-f\text{ in }B_{1},\end{cases}$
we have that
$\sup_{B_{1/2}}u\leq C\left(\inf_{B_{1/2}}u+\|f\|_{L^{n}(B_{1})}\right).$
{exercise}
Show that for $[\lambda,\Lambda]\subseteq(0,\infty)$ there exists some $C\geq 1$ such that the following holds: Let $F\in C({\mathbb{R}}^{n\times n}_{\text{sym}})$ convex, uniformly elliptic with respect to $[\lambda,\Lambda]\subseteq(0,\infty)$, and with $F(0)=0$. Let $u\in C^{4}(B_{1})$ a solution of $F(D^{2}u)=0$ in $B_{1}$, $e\in\partial B_{1}$, and $v_{2}=\partial_{e}^{2}u$. Then
$\max_{\overline{B_{1/2}}}(v_{2})_{-}\leq C\operatorname*{osc}_{B_{1}}u.$
This is effectively a bound on the negative eigenvalues of the Hessian of $u$. Given that $u$ also satisfies $\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u)\leq 0$, the bound on the negatives eigenvalues also implies a similar bound on the positive ones. All in all, we get the $C^{1,1}$ estimate
$\|D^{2}u\|_{L^{\infty}(B_{1/2})}\leq C\operatorname*{osc}_{B_{1}}u.$
Let $F\in C({\mathbb{R}}^{n\times n}_{\text{sym}})$ convex and uniformly elliptic with respect to $[\lambda,\Lambda]\subseteq(0,\infty)$. Given $A\in{\mathbb{R}}^{n\times n}_{\text{sym}}$ and $u\in C^{4}(B_{1})$ a solution of $F(D^{2}u)=0$ in $B_{1}$, consider $v_{A}:=\operatorname{div}(AA^{T}Du)=(AA^{T}):D^{2}u$. Show that $v_{A}$ satisfies $\mathcal{P}_{\lambda,\Lambda}^{-}(D^{2}v_{A})\leq 0$ in $B_{1}$.
Show that for $[\lambda,\Lambda]\subseteq(0,\infty)$ there exists some $\theta\in(0,1)$ such that the following holds: Let $F\in C({\mathbb{R}}^{n\times n}_{\text{sym}})$ convex, uniformly elliptic with respect to $[\lambda,\Lambda]\subseteq(0,\infty)$, and with $F(0)=0$. Let $u\in C^{4}(B_{1})$ a solution of $F(D^{2}u)=0$ in $B_{1}$, with $D^{2}u(0)=0$. Then
$\min_{\begin{subarray}{c}P\in{\mathbb{R}}^{n\times n}_{\text{sym}}\\ P^{2}=P\end{subarray}}v_{P}\geq-1\text{ in }B_{1}\qquad\Rightarrow\qquad\min_{\begin{subarray}{c}P\in{\mathbb{R}}^{n\times n}_{\text{sym}}\\ P^{2}=P\end{subarray}}v_{P}\geq-(1-\theta)\text{ in }B_{1/2}.$
This is effectively a Hölder modulus of continuity for $|(D^{2}u)_{-}|_{\text{op}}$ at the origin. Once again, given that $u$ also satisfies $\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u)\leq 0$, we also get a Hölder modulus of continuity for $|(D^{2}u)_{+}|_{\text{op}}$ at the origin. All in all, we conclude the $C^{2,\alpha}$ estimate
$[D^{2}u]_{C^{0,\alpha}(B_{1/2})}\leq C\operatorname*{osc}_{B_{1}}u.$
4. Degenerate Problems
Some natural questions one may ask is when it is possible to extend the theory whenever uniform ellipticity is not available. This actually the case of some of some of the most important models in elliptic equations such as the minimal surface equation
(4.9) $\operatorname{div}\left(\frac{Du}{\sqrt{1+|Du|^{2}}}\right)=0,$
the $p$-Laplace equation ($p>1$)
$\operatorname{div}(|Du|^{p-2}Du)=0,$
the Monge-Ampère equation ($D^{2}u\geq 0$)
$\det(D^{2}u)=f,$
and in the dynamic case, the porous media equation ($u\geq 0$)
$\partial_{t}u=u\Delta u+|Du|^{2}.$
Free boundary problems such as the obstacle problem
$\min(\Delta u,\varphi-u)=0,$
are also considered degenerate elliptic, given that the uniform ellipticity only has effect outside of the contact region $\{u=\varphi\}$.
In all these cases we observe that $F$ is still monotone increasing on the Hessian, however its rate of growth may go to zero or infinity depending on other features of the solution. Nonetheless, in all of the previous models we find out some additional mechanism which complements the uniform ellipticity in order to establish suitable regularity estimates.
The result we would like to present in this final section considers the case where uniform ellipticity holds if the gradient is sufficiently large.
4.1. Elliptic Equations That Hold Only Where the Gradient is Large
This degeneracy consideration splits the domain of the equation in two regions. One where the gradient is bounded and the solution is automatically Lipschitz, and another where the solution satisfies a uniformly elliptic equation which should enforce some regularity. The problem is that we do not have any control on the boundary between these two regions where the regularity may degenerate.
The following theorem was originally proven in [IS16].
Theorem 4.1.
Given $[\lambda,\Lambda]\subseteq(0,\infty)$ and $\gamma\geq 0$, there exist $\alpha\in(0,1)$ and $C\geq 1$, such that the following holds: Let $u\in C^{1}(B_{1})\cap C^{2}(\{|Du|>\gamma\}\cap B_{1})$, and $f\in C(B_{1})$ non-negative such that
$\begin{cases}\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u)\leq f\text{ in }\{|Du|>\gamma\}\cap B_{1},\\ \mathcal{P}^{+}_{\lambda,\Lambda}(D^{2}u)\geq-f\text{ in }\{|Du|>\gamma\}\cap B_{1},\end{cases}$
Then
$\sup_{r\in(0,1)}r^{-\alpha}\operatorname*{osc}_{B_{r}}u\leq C\left(\operatorname*{osc}_{B_{1}}u+\|f\|_{L^{n}(\{|Du|>\gamma\}\cap B_{1})}\right).$
The family of equations given in the previous theorem are invariant by dilations. When we consider $u_{r}(x):=r^{-\alpha}u(rx)$, $\gamma_{r}:=r^{1-\alpha}\gamma$, and $f_{r}(x):=r^{2-\alpha}f(rx)$, we obtain that
$\begin{cases}\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u_{r})\leq f_{r}\text{ in }\{|Du_{r}|>\gamma_{r}\}\cap B_{1/r},\\ \mathcal{P}^{+}_{\lambda,\Lambda}(D^{2}u_{r})\geq-f_{r}\text{ in }\{|Du_{r}|>\gamma_{r}\}\cap B_{1/r},\\ \|f_{r}\|_{L^{n}(\{|Du_{r}|>\gamma_{r}\}\cap B_{1/r})}=r^{1-\alpha}\|f\|_{L^{n}(\{|Du|>\gamma\}\cap B_{1})}.\end{cases}$
For $\alpha,r\in[0,1]$ the dilations make the threshold $\gamma_{r}$ and the forcing term $f_{r}$ smaller than the original ones.
This means that the proof of such theorem ultimately relies in the proof of a measure estimate, as the one appearing in the Corollary 3.11, which followed from Lemma 3.8 and Lemma 3.10. The remaining steps towards the interior Hölder estimate can be proved with the same arguments we have already presented (diminish of distribution and diminish of oscillation).
Notice as well that, by performing a sufficiently large dilation, we can assume that $\gamma$ is as small as needed. In Lemma 3.10 we observe that the test functions have gradients universally bounded away from zero. Hence the exact same proof applies under a smallness condition on $\gamma$. On the other hand, the result analogous to Lemma 3.8 requires a new idea. Let us present the proof given in [Moo15].
Lemma 4.2.
There exists $\gamma_{0}>0$ sufficiently small such that the following holds. Given $[\lambda,\Lambda]\subseteq(0,\infty)$ there exist $\delta,\theta,\eta\in(0,1)$ and $M\geq 1$, such that for $\gamma\in[0,\gamma_{0}]$, $u\in C^{1}(B_{1})\cap C^{2}(\{|Du|>\gamma\}\cap B_{1})$ non negative, with $\|(\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u))_{+}\|_{L^{n}(\{|Du|>\gamma\}\cap B_{1})}\leq\delta$, it holds that
$u(0)\leq\theta\qquad\Rightarrow\qquad|\{u\leq M\}\cap B_{1}|\geq\eta.$
We use test functions of the form $\varphi_{y_{0}}^{M}=-\frac{M}{2}|x-y_{0}|^{2}$ and their corresponding contact set will be denoted by
$A^{M}_{B}=\bigcup_{y_{0}\in B}\operatorname*{argmin}(u-\varphi_{y_{0}}^{M}).$
Notice that for $\mu\in(0,1)$ and $\theta=(1-\mu)^{2}n/2$, we get that the given hypotheses on Lemma 4.2 imply that for some $c\in[0,\theta]$, $\varphi_{0}^{1}+c$ touches $u$ from below at some point in $\overline{B_{1-\mu}}$. We consider first the scenario where the contact point falls in the region where the gradient is large and the uniform ellipticity is available.
Lemma 4.3.
Given $[\lambda,\Lambda]\subseteq(0,\infty)$ and $\mu\in(0,1/2)$ there exist $\delta,\eta\in(0,1)$, such that for $\gamma\in[0,\mu/2]$, $u\in C^{1}(B_{1})\cap C^{2}(\{|Du|>\gamma\}\cap B_{1})$, with $u\geq\varphi_{0}^{1}$ in $B_{1}$, and
$\|(\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u))_{+}\|_{L^{n}(\{|Du|>\gamma\}\cap B_{1})}\leq\delta,$
it holds that for $R:=\overline{B_{1-\mu}}\setminus B_{\mu}$
$\min_{R}(u-\varphi_{0}^{1})=0\qquad\Rightarrow\qquad|\{u\leq 1\}\cap B_{1}|\geq\eta.$
The following proof is closely related with the idea behind the solution of measure estimate for the Hessian (Problem 13 and its solution). See also Figure 26.
Proof of Lemma 4.3.
Let $x_{0}\in A^{1}_{\{0\}}\cap R$ and consider for $y_{0}\in B_{\mu/10}(x_{0})$
$\varphi_{0}^{1}+(\varphi_{y_{0}}^{1}+\mu^{2}/200)=\varphi_{y_{0}/2}^{2}+c_{y_{0}}.$
For $y_{0}\in B_{\mu/10}(x_{0})$ we have that $\varphi_{y_{0}}^{1}(x_{0})+\mu^{2}/200>0$ and $\varphi_{y_{0}}^{1}+\mu^{2}/200<0$ in ${\mathbb{R}}^{n}\setminus B_{\mu/5}(x_{0})$. Therefore
$\emptyset\neq\operatorname*{argmin}(u-\varphi_{y_{0}/2}^{2})\subseteq\{u\leq\mu^{2}/200\}\cap B_{\mu/5}(x_{0})\subseteq B_{1}.$
For every $x_{1}\in\operatorname*{argmin}(u-\varphi_{y_{0}/2}^{2})$ we get that
$|Du(x_{1})|$ $=2|x_{1}-y_{0}/2|$
$\geq 2(|x_{0}-x_{0}/2|-|x_{0}-x_{1}|-|y_{0}/2-x_{0}/2|)$
$\geq 2(|x_{0}|/2-\mu/5-\mu/20)$
$>\mu/2$
$\geq\gamma.$
Finally, by the same argument in the proof of Lemma 3.8, using the uniform ellipticity at the contact points and the area formula, we conclude that
|$|\{u\leq 1\}\cap B_{1}|\geq|A^{2}_{B_{\mu/20}(x_{0}/2)}|\geq\eta.$
Now we consider the alternative scenario when the contact point falls in the region where the gradient is small. The following result could be considered as a localization-type estimate.
Lemma 4.4.
There exist $\mu,r\in(0,1)$ with $r+(1-\mu)/2<1$, such that for $u\in C(B_{1})$ with $u\geq\varphi_{0}^{1}$, and $x_{0}\in\partial B_{r}$ it holds that
$\min_{\overline{B_{\mu}}}(u-\varphi_{0}^{1})=0\qquad\Rightarrow\qquad\emptyset\neq A_{\{x_{0}\}}^{2}\subseteq\{u\leq 1\}\cap\overline{B_{(1-\mu)/2}(x_{0})}.$
Proof of Lemma 4.4.
We will choose $\mu$ and $r$ such that
(4.10) $m:=\max_{\overline{B_{\mu}}}(\varphi_{0}^{1}-\varphi_{x_{0}}^{2})\leq\min_{{\mathbb{R}}^{n}\setminus\overline{B_{(1-\mu)/2}(x_{0})}}(\varphi_{0}^{1}-\varphi_{x_{0}}^{2})\leq 1.$
Once this is done we observe that the paraboloid $\varphi^{2}_{x_{0}}+m$ is greater than $\varphi_{0}^{1}$ over $\overline{B_{\mu}}$, and less than $\varphi_{0}^{1}$ outside $\overline{B_{(1-\mu)/2}(x_{0})}$. Hence, the hypotheses on $u$ imply that $\varphi_{x_{0}}^{2}+m$ must cross $u$ in $\overline{B_{(1-\mu)/2}(x_{0})}$. By lowering the paraboloid we find a contact point in $\overline{B_{(1-\mu)/2}(x_{0})}$. See Figure 21.
Figure 21. We arrange $\mu$ and $r$ and $m$ such that the graph of $u$ enters the region $\{(x,y)\in{\mathbb{R}}^{n}\times{\mathbb{R}}\,:\,\varphi_{0}^{1}(x)\leq y\leq\varphi_{x_{0}}^{2}(x)+m\}\subseteq B_{(1-\mu)/2}(x_{0})\times(-\infty,1]$.
We compute
$\varphi_{0}^{1}(x)-\varphi_{x_{0}}^{2}(x)=\frac{1}{2}|x-2x_{0}|^{2}+C_{r}.$
The maximum of this convex paraboloid over $\overline{B_{\mu}}$ is achieved at the boundary
$\max_{\overline{B_{\mu}}}(\varphi_{0}^{1}-\varphi_{x_{0}}^{2})=\frac{1}{2}(\mu+2r)^{2}+C_{r}.$
If we set $r\leq(1-\mu)/2$, so that $2x_{0}$, the vertex of $(\varphi_{0}^{1}-\varphi_{x_{0}}^{2})$, falls in $\overline{B_{(1-\mu)/2}(x_{0})}$, we observe that the minimum of $(\varphi_{0}^{1}-\varphi_{x_{0}}^{2})$ over the complement of $\overline{B_{(1-\mu)/2}(x_{0})}$ is also attained over the boundary and in particular it is less than 1. Moreover,
$\min_{{\mathbb{R}}^{n}\setminus\overline{B_{(1-\mu)/2}(x_{0})}}(\varphi_{0}^{1}-\varphi_{x_{0}}^{2})=\frac{1}{2}\left(\frac{1-\mu}{2}-r\right)^{2}+C_{r}.$
Then we need $\mu+2r\leq\frac{1-\mu}{2}-r$ to have (4.10). Putting all the restrictions together, we find out that the result holds for any $\mu\in(0,1/3)$ and $r\in(0,1/6-\mu/2)$. ∎
Proof of Lemma 4.2.
Let us show the result instead in $B_{3\sqrt{n}}$ and under the same hypotheses. We rescale the results in Lemma 4.3 and Lemma 4.4 as well with $\mu=1/6$ and $r=1/12$.
Let
$m_{k}(x_{0}):=\inf_{B_{3\sqrt{n}}}(u-\varphi_{x_{0}}^{2^{k}}),\qquad R_{r}(x_{0}):=\overline{B_{5r\sqrt{n}/2}(x_{0})}\setminus B_{r\sqrt{n}/2}(x_{0}).$
By choosing $\theta:=(1-\mu)^{2}9n/2=225n/72$, we get that $m_{0}(0)\leq 225n/72$ and
$\min_{\overline{B_{5\sqrt{n}/2}}}(u-(\varphi_{0}^{1}+m_{0}(0)))=0.$
We select a set of dyadic cubes according to the following algorithm: Initially we consider the zeroth generation of dyadic cubes $\{Q_{1}\}$. In general, for the $k^{th}$ stage ($k\geq 0$), we have a subset of cubes $\{Q_{1/2^{k}}(x_{0})\}$ from the $k^{th}$ generation of dyadic cubes.
The cube $Q_{1/2^{k}}(x_{0})$ is selected if
$\min_{R_{1/2^{k}}(x_{0})}(u-(\varphi_{x_{0}}^{2^{k}}+m_{k}(x_{0})))=0.$
Otherwise, the descendants of $Q_{1/2^{k}}(x_{0})$ will be considered in the next stage.
By an inductive argument using Lemma 4.4, we obtain that for every cube $Q_{1/2^{k}}(x_{0})$ that is considered in the $k^{th}$ iteration we also have that
$A^{2^{k}}_{\{x_{0}\}}\subseteq\{u\leq m_{k}(x_{0})\}\cap\overline{B_{5\sqrt{n}/2^{k+1}}(x_{0})}.$
Also from Lemma 4.4 we get that if $Q_{1/2^{k+1}}(x_{1})$ is a cube considered in the $(k+1)^{th}$ iteration, descendant of $Q_{1/2^{k}}(x_{0})$, then
$m_{k+1}(x_{1})\leq m_{k}(x_{0})+2^{-k}\leq 225n/72+2.$
The last inequality follows by induction.
By Lemma 4.3 we get that for the set of selected cubes and $M:=225n/72+3$,
|$|\{u\leq M\}\cap B_{3\sqrt{n}/2^{k}}(x_{0})|\geq\eta_{0}|B_{3\sqrt{n}/2^{k}}(x_{0})|.$
If $x_{0}\in Q_{1}$ is not covered by the selected cubes we find a nested sequence of dyadic cubes $Q_{1}\supseteq Q_{1/2}(x_{1})\supseteq\ldots\ni x_{0}$ such that
$\min_{\overline{B_{\sqrt{n}/2^{k+1}}(x_{k})}}(u-(\varphi_{x_{k}}^{2^{k}}+m_{k}(x_{k})))=0.$
Which implies $u(x_{0})\leq M$.
By combining the two alternatives we recover the desired estimate on the measure of $\{u\leq M\}\cap B_{3\sqrt{n}}$. If the measure of the set of points in $Q_{1}$ not covered is at least $1/2$, the result would follow for any $\eta\leq 1/2$. Otherwise, we use Vitali’s covering lemma to extract a disjoint sequence of balls covering a set of measure at least $25^{-n}/$ 2, where each ball has
|$|\{u\leq M\}\cap B|\geq\eta_{0}|B|.$
Adding these measures we recover the desired result for $\eta:=5^{-n}\eta_{0}/2$. ∎
{exercise}
Assume the hypotheses of Lemma 4.2 with $\theta=1$. Consider the family of test functions of the form
$\varphi_{y_{0}}(x):=-C_{0}|x-y_{0}|^{1/2},$
and the contact set
$A:=\bigcup_{y_{0}\in\{u>M\}\cap B_{1/4}}\operatorname*{argmin}(u-\varphi_{y_{0}}).$
Show that for some appropriated choice of $C_{0},M\geq 1$ and $\gamma_{0},\delta,\eta\in(0,1)$, we have that $A\subseteq\{u\leq M\}$ with $|A|\geq\eta$.
4.1.1. Further Developments
The methods described in [IS16, Moo15] have proven to be remarkably versatile, extending their applicability to various equations. The cusps test functions introduced in [IS16] were used by Silvestre and Schwab to extend the interior regularity estimates for parabolic integro-differential equations in [SS16]. Recently, Pimentel, Santos, and Teixeira further advanced this idea to derive higher-order fractional estimates in [PST22]. Moreover, in collaboration with Santos, we revisited the regularity theory for the porous medium equation, as presented in [CLS23], by adapting Mooney’s algorithm to a parabolic context.
4.2. Quasi-Harnack Inequality
Degeneracy can also manifest itself across scales. For example, when modeling a PDE using finite difference schemes, the continuous formulation of uniform ellipticity breaks down at the level of the discretization. However, if the numerical scheme approximates a uniformly elliptic equation, we expect that the discrete solution will approximate the continuous solution over large scales, inheriting with it the classical manifestations of uniform ellipticity [KT96].
A recent work by De Silva and Savin in [DSS21] proposes a weak notion of solution for equations where the uniform ellipticity manifests from a given scale onward. Their main result is a weak Harnack type-inequality. In this case, the alternative mechanism that compensates for the lack of uniform ellipticity is a measure estimate at microscopic scales. We discuss this result in more detail in the recent survey [CL23]. It is worth noticing that the proof of their main measure estimate relies once again in a dyadic cube algorithm selection [DSS21, Lemma 3.4].
Besides the already discussed applications to numerical schemes, the quasi-Harnack inequality can also be applied in the homogenization of elliptic problems with degeneracies, as studied in [AS14]. It was also shown in [DSS21] that integro-differential uniformly elliptic equations of order $\sigma$ close to two, also fit in the framework of the quasi-Harnack inequality. In this way, it provides a new proof to the Harnack inequality of Caffarelli and Silvestre [CS09].
A particular attractive feature of the quasi-Harnack inequality is that it does not require an actual partial differential equation. This is the idea behind the previous mentioned applications. This versatility allowed to implement this approach to almost minimizers in calculus of variations in [DSS20].
4.3. Small Perturbation Solutions for Elliptic Equations
The previous framework for degenerate equations falls short when addressing crucial cases in calculus of variations, such as the minimal surface equation (4.9). In this context, the ellipticity degenerates as the gradient grows. The approach to regularity in this scenario significantly differs from the method presented in our current discussion. In this short section we would like to highlight some of the main references and refeer to [CL23] for a more technical overview of the approach.
Caffarelli introduced a perturbative method in [Caf89] for establishing higher regularity estimates for solutions of uniformly elliptic equations, often referred to as regularity by compactness or the improvement of flatness. Here, “flatness” means that a solution is assumed to be uniformly close to a prescribed profile. This innovative strategy was inspired by De Giorgi’s regularity theorem for minimal surfaces [DG60]. The core idea is that if a solution closely resembles a smooth solution, it inherits the estimates of the corresponding linearization.
This approach swiftly provided alternative proofs for regularity estimates in certain degenerate equations. Caffarelli and Córdoba applied it to the minimal surface equation in [CC93], while Wang explored estimates for the $p$-Laplace equation in [Wan94].
In [Sav07], Savin extended these estimates to encompass operators $F=F(M,p,z,x)$ that only need to be uniformly elliptic in the vicinity of a smooth solutions (taken as zero in [Sav07]). Notably, this result enabled the treatment of equations that become degenerate as $(M,p,z)$ grows large, complementing the lack of uniform ellipticity whenever $(M,p,z)$ belongs to an unbounded regions. This approach allowed Savin to establish a long standing conjecture of De Giorgi about the level sets of semi-linear equations arising from the Ginzburg-Landau energy in [Sav09].
4.3.1. Further Developments
The groundbreaking work in [Sav07] has paved the way for multiple significant developments in the field. While there are too numerous to comprehensively cover all of them here, some noteworthy results include the following main research directions:
Free Boundary Problems: De Silva developed the perturbative approach for the Bernoulli free boundary problem (also known as the One-Phase or Alt-Caffarelli problem), starting in [DS11]. In this case, the free boundary condition linearizes to a zero Neumann type problem on the half-space, or equivalently the non-local problem $\Delta^{1/2}u=0$. A parabolic counterpart was studied in [CLGS19]. A setup leading to the obstacle problem for $\Delta^{1/2}$ was studied in [CLS19] and used in [DPSV21] to analyze the branching points in a two phase free boundary problem.
Partial and Boundary Regularity for Uniformly Elliptic Equations: Armstrong, Silvestre, and Smart applied this method to develop partial regularity results for fully non-linear equations in [ASS12]. One should keep in mind that in general, uniformly elliptic problems may develop singularities as the one exhibited by [NV13]. However, the interesting result proved in [ASS12], shows that those singularities must have positive Hausdorff co-dimension (and hence zero measure). Silvestre and Sirakov proved in [SS14] that the singular set can only be found in the interior of the domain of the equation, once again relying on [Sav07].
Optimal Regularity for Degenerate Equations: Colombo and Figalli developed regularity estimates for degenerate equations arising from traffic congestion models in [CF14]. In collaboration with Pimentel, we demonstrated in [CLP21] the continuity of $|Du|$, where $u$ solves the gradient-constrained problem $\max(1-|Du|,\Delta u+1)=0$.
In [Le18], the paraboloid method is implemented to show a Harnack inequality for singular elliptic equations which are $(D^{2}u)^{-1}$-like. These arise as the linearizations of the Monge-Ampère equation.
Non-Local Minimal Surfaces: Caffarelli, Roquejoffre, and Savin established regularity estimates for non-local minimal surfaces in [CRS10].
Parabolic and Non-Local Counterparts: The technique was extended to parabolic problems by Wang in [Wan13], and for non-local equation by Yu in [Yu17].
5. Open Problems
One of the most famous open problems in the field is to characterize the dimension for which the translation invariant equation $F(D^{2}u)=0$ has $C^{2}$ estimates.
It was shown by Nirenberg using complex analysis tools in [Nir53], that $C^{2,\alpha}$ estimates are available in dimension 2. Recall that in Problem 20 we saw that $C^{2,\alpha}$ estimates are available for convex/concave operators in any dimension. Some further generalizations have been studied in [CC03, CY00]. However, it is known that in general we may find singularities starting in dimension $5$, [NV13].
Do solutions of the uniformly elliptic equation $F(D^{2}u)=0$ have interior $C^{2}$ estimates or do singularities may also develop in dimension 3 or 4?
Whenever singularities develop for the equation $F(D^{2}u)=0$, it was shown in [ASS12] that they have positive Hausdorff co-dimension. This co-dimension is at least the exponent $\varepsilon$ in the measure estimate for the Hessian (Problem 15), which yields the question of how large could $\varepsilon$ actually be. It was observed in the same article that the in dimension 2, one must have that $\varepsilon\leq C(\Lambda/\lambda)^{-1}$ and conjectured that this is actually optimal.
In a recent approach to the Krylov-Safonov Theory by Mooney [Moo19], it has been shown that one can take $\varepsilon\geq C_{n}(\Lambda/\lambda)^{1-n}$, settling the previous conjecture in dimension 2; see also [Le20]. It was conjectured in [Moo19] that this estimate is actually the optimal one in any dimension, however its verification remains open.
One may also wonder about the optimal integrability exponent $p$ allowed as a forcing term in a uniformly elliptic equation in order to get interior Hölder estimates. As we know from the Sobolev and Morrey embeddings, this exponent should be strictly greater than $n/2$. For uniformly elliptic operators, we just saw that the theory holds for $p\geq n$. Due to fundamental estimates arising from harmonic analysis in [FS84], it was shown in [Esc93] that the optimal $p$ can always be taken slightly smaller in the interval $(n/2,n)$. The actual dependence of the optimal integrability exponent $p$ on the dimension and the ellipticity constants remains open.
A quite challenging question related to the integrability exponent for non-local equations of order $2\sigma\in(0,2)$ is the following one: In contrast to the second order theory, the non-local ABP estimate proved in [CS09] depends on the $L^{\infty}$-norm for the forcing term. It is expected that a maximum principle also holds in terms of the $L^{p}$-norm for the forcing term, for some $p\in(n/(2\sigma),\infty)$. For the fractional Laplacians this is a consequence of the $L^{p}$ bounds for the Riesz transform and the Sobolev embedding. In the uniformly elliptic setting, the proof of this conjecture is only known under restrictive assumptions on the kernel [GS12, Kit20].
Similar questions can also be posed regarding the optimal Hölder exponent. So far, it seems that new techniques may be required in dimension greater or equal than 3. See [Tei20] for some conjectures in the case of the $p$-Laplacian.
The article [IS16], which studied Hölder estimates for elliptic equations that hold whenever the gradient is large (Theorem 4.1), posed two interesting problems. The first one was to extend the estimate for equations with drift term. This was solved in [Moo15], whose strategy was presented in this note (in the absence of drift).
The second problem remains open: Is it possible to get a Hölder estimate on a solution of the parabolic problem
$\begin{cases}\partial_{t}u-\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u)\geq-|f|\text{ in }\{|Du|>\gamma\}\cap B_{1}\times(-1,0],\\ \partial_{t}u-\mathcal{P}^{+}_{\lambda,\Lambda}(D^{2}u)\leq|f|\text{ in }\{|Du|>\gamma\}\cap B_{1}\times(-1,0].\end{cases}$
Notice that any $u=u(t)$, independent of $x$, is automatically a solution for any forcing term $f$. This shows that we can not expect a regularity estimate in time. Perhaps a Hölder estimate exclusively in space should be true.
Other reasonable extension in the static case would be to replace the constant $\gamma$ by some non-negative function $\gamma\in L^{p}$. For $p>n$, we have by Morrey’s inequality that $|Du|\leq\gamma$ implies $u\in C^{0,1-n/p}$. Would the Theorem 4.1 hold under the assumption $\gamma\in L^{p}$ for some $p>n$?
6. Ideas and References for the Problems
1. Use the fundamental theorem of calculus and Hölder’s inequality.
2. Let $\rho\in(0,r)$ and $\theta\in\partial B_{1}$. By the fundamental theorem of calculus
[…]
Integrating over $\theta\in\partial B_{1}$
[…]
A final integration from $\rho=0$ to $\rho=r$, after having multiplied by $\rho^{n-1}$, gives the desired identity.
3. By the triangle inequality $\operatorname*{osc}_{B_{r}}u\leq 2\sup_{B_{r}}|u-u(0)|$. Once again by the triangle inequality, we get that for $x\in B_{r}$ with $r\in(0,1/2)$
$|u(x)-u(0)|\leq\fint_{B_{r}}|u-u(0)|+\fint_{B_{r}(x)}|u-u(x)|.$
Hence the desired result follows by Hölder’s inequality using that $1/|x|^{n-1}\in L^{q}(B_{2})$ for $q>n/(n-1)$.
1. Use that for $r\in(0,1/3)$ and $x,y\in B_{r}$
$B_{1-3r}(y)\subseteq B_{1-r}(x)\subseteq B_{1}.$
2. Follows by using $u(0)$ instead of $\fint_{B_{1/3}}u$ in the proof of Lemma 2.6. We also replace $B_{1/3}$ by $B_{1/4}$.
1. Consider an upper supporting plane for $u$ at $x_{0}\in\operatorname*{argmin}_{\overline{B_{r}}}u\subseteq\partial B_{r}$.
2. The set $\{u\geq\sup_{B_{r}}u\}$ must contain a ball of diameter $(1-r)$.
1. Use polar coordinates for the integration defininf the convolution.
2. Using the Taylor expansion around zero for $x=\varepsilon\theta$ and $\theta\in B_{1}$
$u(x)-u(0)=\varepsilon Du(0)\cdot\theta+\frac{\varepsilon^{2}}{2}\theta\cdot D^{2}u(0)\theta+o(\varepsilon^{2}).$
The first term is odd and its integral over $B_{\varepsilon}$ vanishes. The last term is an error that vanishes faster than $\varepsilon^{2}$ as $\varepsilon\to 0$. Hence we just keep the quadratic terms in the limit. Once again, we notice the odd symmetry of $\theta_{i}\theta_{j}$ for $i\neq j$ to keep just the quadratic terms of the form $\theta_{i}^{2}$ in the integration.
By using the spherical symmetry once again
[…]
Details can be consulted in [HL11, Theorem 1.8 in Section 1.2].
1. Show first that the following variation of the mean value property for $v=\min(u,0)$
$\fint_{B_{1}}v\leq v(0)+C\|(\Delta u)_{+}\|_{L^{p}(\{u<0\}\cap B_{1})}.$
Then the Hölder modulus of continuity will follow from the diminish of oscillation approach, just like it was done for the interior Hölder estimate.
For $\varepsilon>0$, consider $\phi_{\varepsilon}\in C^{\infty}({\mathbb{R}})$, non-decreasing and concave, such that
$\phi_{\varepsilon}(z)=\begin{cases}-z\text{ if }z\leq-\varepsilon,\\ 0\text{ if }z\geq 0.\end{cases}$
Then we apply the mean value property to $v_{\varepsilon}:=\varphi_{\varepsilon}(u)$ and conclude after sending $\varepsilon\to 0^{+}$. Notice that
$(\Delta v_{\varepsilon})_{+}=(\phi_{\varepsilon}^{\prime}\Delta u+\phi_{\varepsilon}^{\prime\prime}|Du|^{2})_{+}\leq\mathbbm{1}_{\{u<0\}}(\Delta u)_{+}.$
1. Let $\theta$ to be fixed sufficiently small and assume by contradiction that $u<(1+\theta)u(0)$ in $B_{1}$. The idea is that $\|u_{+}\|_{L^{\varepsilon}(B_{1})}\leq 1$ and the mean value property for $-u$ compete in opposite directions when we estimate the measure of $\{u\geq u(0)/2\}\cap B_{1/3}$.
Let $M\geq 1$ be sufficiently large such that by Markov’s inequality
(6.11) |$\frac{|\{u\geq u(0)/2\}\cap B_{1/3}|}{|B_{1/3}|}\leq\frac{(u(0)/2)^{-\varepsilon}}{|B_{1/3}|}\|u_{+}\|_{L^{\varepsilon}(B_{1})}^{\varepsilon}\leq CM^{-\varepsilon}<\frac{1}{2}.$
By a combination of the Markov inequality and the mean value property applied to $v=(1+\theta)u(0)-u>0$
|$\frac{|\{u<u(0)/2\}\cap B_{1/3}|}{|B_{1/3}|}$ |$\leq\frac{|\{v>u(0)/2\}\cap B_{1/3}|}{|B_{1/3}|}$
$\leq 2u(0)^{-1}\fint_{B_{1/3}}v$
$\leq 2u(0)^{-1}(\theta u(0)+C),$
$\leq 2\theta+CM^{-1}.$
By taking $M$ sufficiently large $\theta$ sufficiently small we enforce the right-hand side to be less than $1/2$, contradicting (6.11).
2. The idea proceeds once again by contradiction. We will use the previous part to obtain a sequence of points $\{x_{k}\}\subseteq B_{3/4}$ where $u(x_{k})\geq(1+\theta)u(x_{k-1})\geq\ldots\geq(1+\theta)^{k}M\to\infty$, contradicting the continuity of $u$. To prevent that the points escape from $B_{3/4}$, we will also show that we can make this choice with $|x_{k}-x_{k-1}|$ controlled by a geometric vanishing sequence.
The same argument as in the previous part shows that for some $\sigma\geq 1$ sufficiently large, it holds that if $B_{3r}(x_{0})\subseteq B_{1}$ with $r=\sigma u(x_{0})^{-\varepsilon/n}$, then
$\begin{cases}\|u_{+}\|_{L^{\varepsilon}(B_{1})}\leq 1,\\ \|(\Delta u)_{-}\|_{L^{p}(B_{1})}\leq 1,\\ u(x_{0})\geq M,\end{cases}\qquad\Rightarrow\qquad\sup_{B_{r}(x_{0})}u\geq(1+\theta)u(x_{0}).$
Indeed, we only need to modify the upper estimate on the measure of $\{u\geq u(x_{0})/2\}\cap B_{r}(x_{0})$ by
|$\frac{|\{u\geq u(x_{0})/2\}\cap B_{r}(x_{0})|}{|B_{r}(x_{0})|}\leq C\sigma^{-n}<\frac{1}{2}.$
By choosing $M$ even larger we can enforce that
$\sigma M^{-\varepsilon/n}\sum_{k=0}^{\infty}(1+\theta)^{-k\varepsilon/n}\leq\frac{1}{4}.$
Then, if we assume that $u(x_{0})\geq M$ for some $x_{0}\in B_{1/2}$ we can then construct a sequence of points $x_{1},x_{2},\ldots$ inductively such that
$x_{k+1}\in\operatorname*{argmax}_{\overline{B_{r_{k}}(x_{k})}}u,\qquad r_{k}:=\sigma u(x_{k})^{-\varepsilon/n}.$
This is the desired sequence which leads to the expected contradiction.
Details can be consulted in [CC95, Lemma 4.7 and the proof of Lemma 4.4].
1. (Mean Value Property) The main idea is to consider $v=\eta u$, for some non-negative and compactly supported cut-off function $\eta\in C^{\infty}_{c}(B_{1}\times(-1,0])$, so that $v$ would now satisfy the heat equation in ${\mathbb{R}}^{n}\times(-\infty,0]$.
Given that
$\partial_{t}v-\Delta v=\eta f+(\partial_{t}\eta-\Delta\eta)u-2D\eta\cdot Du.$
Then, if $\eta(0,0)=1$,
[…]
Hence we take
$\begin{cases}K_{2}=\eta H_{-t},\\ K_{1}=[(\partial_{t}\eta+\Delta\eta)+t^{-1}x\cdot D\eta]H_{-t}.\end{cases}$
We can easily see that $K_{2}\geq 0$, now we need to choose $\eta$ such that $(\partial_{t}\eta+\Delta\eta)+t^{-1}x\cdot D\eta\geq 0$. The function $\eta=(1-2|x|^{2}+4nt)_{+}$ does the work, but it is not smooth and the computations above can not be directly applied. Instead we may consider a modification of this candidate by approximating the positive part with a smooth function as was done for example in the solution of Problem 5.
See [Eva10, Chapter 2.3] for an alternative way to construct the mean value property for the heat equation.
2. (Weak Harnack) In the previous construction we have that for some $\rho\in(0,1/2)$ sufficiently small, $\min_{\overline{B_{\rho}}\times[-\rho,-2\rho]}K_{1}>0$. The singularity of the heat kernel at the origin gives that
$\iint fK_{2}\leq C\|f\|_{L^{q}_{t}((-1,0]\to L^{p}_{x}(B_{1}))}.$
Then
$\iint_{B_{\rho}\times(-\rho,-2\rho]}u\leq C(u(0,0)+\|f_{-}\|_{L^{q}_{t}((-1,0]\to L^{p}_{x}(B_{1}))}).$
The conclusion can finally be obtained from this estimate by a geometric argument, similar to the one illustrated in Figure 3 and adapted to the parabolic setting.
1 and 2. Details can be consulted in [SS05, Section 3.3].
Figure 22. For the ABP maximum principle we consider the hyper-planes that support the graph of $u$ from below. Any plane with slope less than $m/2$ can be slid from below until it touches the graph of $u$.
1. The geometric idea is illustrated in Figure 22. For $m=\max_{\overline{B_{1}}}u_{-}$ and $T=Du\colon A\to{\mathbb{R}}^{n}$ we get that $B_{m/2}\subseteq T(A)$. The Jacobian of $T$ for $x_{0}\in A$ can be estimated using that if $e\in\operatorname{eig}(D^{2}u(x_{0}))$ then $0\leq\lambda e\leq\mathcal{P}_{\lambda,\Lambda}^{-}(D^{2}u(x_{0}))$. Therefore $0\leq\det(DT(x_{0}))\leq C(\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u(x_{0})))^{n}$.
Details can be consulted in [CC95, Chapter 3].
2. Consider the mapping as above using the set of slopes in the convex envelope of
$B_{|u(x_{0})|/\operatorname*{diam}(\Omega)}\cup\{|u(x_{0})|/\operatorname*{dist}(x_{0},\partial\Omega)\}.$
Details can be consulted in [Fig17, Theorem 2.8].
1. Follows directly from the definition.
2. We focus on a compact set $\overline{B_{r}(x_{0})}\subseteq\Omega$ where $u$ is bounded and uniformly continuous. Let us say that $M:=\operatorname*{osc}_{B_{r}(x_{0})}u<\infty$, and $\omega$ is an admissible modulus of continuity for $u$ over $B_{r}(x_{0})$.
Then we have that for $\rho:=\sqrt{2M\varepsilon}<r$ and $y_{0}\in B_{r-\rho}(x_{0})$, the infimum that defines $u_{\varepsilon}$ is actually a minimum that is achieved over $\overline{B_{\rho}(y_{0})}$. Let
$x_{0}\in\operatorname*{argmin}_{\overline{B_{\rho}(y_{0})}}(u-\varphi_{y_{0}}^{1/\varepsilon}).$
Then
$0\leq u(y_{0})-u_{\varepsilon}(y_{0})=u(y_{0})-u(x_{0})-\frac{1}{2\varepsilon}|x_{0}-y_{0}|^{2}\leq u(y_{0})-u(x_{0})\leq\omega(\rho)=\omega(\sqrt{2M\varepsilon}).$
3. One can notice that $x\mapsto u_{\varepsilon}(x)+|x|^{2}/(2\varepsilon)$ is the infimum of affine functions. Alternatively one can show that every point in the graph of $u_{\varepsilon}$ has a supporting paraboloid of the form $\varphi_{x_{0}}^{-1/\varepsilon}+C$ from above, see Figure 23.
4. Follows from a direct computation.
5. Prove that for every $y_{0}\in{\mathbb{R}}^{n}$, $(\varphi_{y_{0}}^{1/\varepsilon}+u_{\varepsilon}(y_{0}))$ is the largest paraboloid centered at $y_{0}$, and with Hessian $-\varepsilon^{-1}I$, that remains below $u$ in $\Omega$.
1. and 2. (Measure Estimate and Localization with Small Drift) Notice that in the proof of Lemma 3.8 and Lemma 3.10 we get that the contact set is contained in a set where $|Du|$ is universally bounded. Hence, the drift term can be absorbed once we assume that $\|b\|_{L^{n}(B_{1})}$ is sufficiently small.
Figure 23. The paraboloid $\varphi_{x_{0}}^{-1/\varepsilon}$ must be tangent to $u_{\varepsilon}$ from above. Otherwise we would be able to find at least one vertex of the form $(y_{1},u_{\varepsilon}(y_{1}))$ strictly on top of the graph of $\varphi_{x_{0}}^{-1/\varepsilon}$. However, by symmetry of the paraboloids we observe that in such case the paraboloid $\varphi_{y_{1}}^{1/\varepsilon}+u_{\varepsilon}(y_{1})$ must be larger than $u$ at $x_{0}$, which contradicts the definition of $u_{\varepsilon}(y_{1})$.
3. (ABP Maximum Principle) Let $m=\max_{\overline{B_{1}}}u_{-}$ and consider for $r\in(0,1/2]$, the contact sets
$A_{r}:=\bigcup_{p\in B_{rm}\setminus B_{rm/2}}\operatorname*{argmin}(u-\varphi_{p}).$
Show then that for some $C>1$, independent of $r$,
$1\leq C(\|b\|^{n}_{L^{n}(A_{r})}+(rm)^{-n}\|f\|_{L^{n}(A_{r})}^{n}).$
Add then the results for $r=1/2,1/4,\ldots,1/2^{k}$ such that $k=\lfloor 2C\Lambda^{n}\rfloor+1$, and use that $\|b\|_{L^{n}(B_{1})}\leq\Lambda$.
The proof in [HL11, Theorem 2.21] also relies on giving different weights to the contact sets $A_{r}$
4. (Localized Measure Estimate) The idea is to localize the minimum to some $B_{\varepsilon}(x_{0})\subseteq B_{\rho}$ such that $\|b\|_{L^{n}(B_{\varepsilon}(x_{0}))}$ is sufficiently small and $\min_{\overline{B_{\varepsilon}(x_{0})}}u$ is not too large, both quantities measured in terms of the universal constants. After having done this, the result in part 1 concludes the argument.
The condition that $\|b\|_{L^{n}(B_{\varepsilon}(x_{0}))}$ is small can be arranged from a geometric argument. The maximum principle allows to find an annular region $B_{r+\varepsilon}\setminus B_{r-\varepsilon}\subseteq B_{1}\setminus B_{1/2}$ such that $\|b\|_{L^{n}(B_{r+\varepsilon}\setminus B_{r-\varepsilon})}$ is sufficiently small and $\min_{\partial B_{r}}u\leq 2$. Finally $B_{\varepsilon}(x_{0})$ and $B_{r+\varepsilon}\setminus B_{r-\varepsilon}$ can be connected by some thin cylinder $T$ where $\|b\|_{L^{n}(T)}$ is also small. This geometric arrangement allows to transmit the minimum around $\partial B_{r}$ to $\overline{B_{\varepsilon}(x_{0})}$ using the localization lemma with small drift.
Details can be consulted in [Moo15, Proposition 6.1].
Figure 24. The test functions are planes of the form $x\mapsto p\cdot x-h$ that at time $t=-1$ are below $u$. We bring them form the past until they touch the graph of $u$ at some time $t_{0}$ and some point $x_{0}$.
1. (ABP Maximum Principle) The test functions and contact set are illustrated in the Figure 24. Let $T\colon A\to{\mathbb{R}}^{n}\times{\mathbb{R}}$ such that $(p_{0},h_{0})=T(x_{0},t_{0})$ if and only if $u>\varphi_{p_{0},h_{0}}$ in $B_{1}\times(-1,t_{0})$, with equality at $(x_{0},t_{0})$. This means that $T=(Du,x\cdot Du-u)$.
For $m=\max_{\overline{B_{1}}}u_{-}$ we get that
$\{(p,h)\in{\mathbb{R}}^{n}\times{\mathbb{R}}\,:\,h\in(0,m/2),p\in B_{h/2}\}\subseteq T(A).$
The Jacobian of $T$ is computed by
$\det\begin{pmatrix}D^{2}u&D(\partial_{t}u)\\ x^{T}D^{2}u&x\cdot D(\partial_{t}u)-\partial_{t}u\end{pmatrix}=-\partial_{t}u\det(D^{2}u).$
For $(x_{0},t_{0})\in A$ and $e\in\operatorname{eig}(D^{2}u(x_{0},t_{0}))$
$0\leq\lambda e\leq\partial_{t}u(x_{0},t_{0})+f(x_{0},t_{0})\leq f(x_{0},t_{0}).$
Therefore $0\leq\det(D_{x,t}T)\leq Cf^{n+1}$ over the contact set.
Details can be consulted in [Tso85].
2. (Measure Estimate) Let $\varphi_{y_{0},s_{0}}(x,t):=(t-s_{0})-2|x-y_{0}|^{2}$ for $(y_{0},s_{0})\in{\mathbb{R}}^{n}\times{\mathbb{R}}$. Notice that $\varphi_{0,-1}(0,0)>0$ and $\varphi_{0,-1}<0$ in $\partial_{\text{par}}(B_{1}\times(-1,0])$. For some $\rho\in(0,1)$ sufficiently small and $(y_{0},s_{0})\in B:=B_{\rho}\times[-1,-1+\rho)$, we still have that $\varphi_{y_{0},s_{0}}<0$ in $\partial_{\text{par}}(B_{1}\times(-1,0])$ and $\theta:=\min_{(y_{0},s_{0})\in B}\varphi_{y_{0},s_{0}}(0,0)>0$.
If $u$ is non-negative and $u(0,0)\leq\theta$, then for every $(y_{0},s_{0})\in B$, the graph of $\varphi_{y_{0},s_{0}}$ must eventually touch the graph of $u$ from below at some point in $B_{1}\times(-1,0]$. Consider in this way the contact set
$A:=\{(x_{0},t_{0})\in B_{1}\times(-1,0]\,:\,\text{For some $(y_{0},s_{0})\in B$,}$
$u>\varphi_{y_{0},s_{0}}$ in $B_{1}\times(-1,t_{0})$,
$\text{and $u(x_{0},t_{0})=\varphi_{y_{0},s_{0}}(x_{0},t_{0})$}\},$
and the surjective map $T\colon A\to B$ such that $(y_{0},s_{0})=T(x_{0},t_{0})$ if and only if $u>\varphi_{y_{0},s_{0}}$ in $B_{1}\times(-1,t_{0})$, with equality at $(x_{0},t_{0})$. This means that
$T=\left(x+\frac{1}{4}Du,t-u-\frac{1}{8}|Du|^{2}\right),$
and its Jacobian is
$\det(D_{x,t}T)=\left(1-\partial_{t}u\right)\det\left(I+\frac{1}{4}D^{2}u\right).$
Using the uniform ellipticity we get that
$0\leq\det(D_{x,t}T)\leq C(1+f^{n+1}).$
3. (Localization) The argument is an adaptation of the Lemma 3.10. It now requires to design a test function $\varphi=\varphi(x,t)\colon B_{1+\rho/2}\times(-1-\rho/2,0]\to{\mathbb{R}}$ such that
$\begin{cases}\varphi<0\text{ in }B_{1+\rho/2}\times(-1-\rho/2,0]\setminus B_{1-\rho/2}\times(-1,0],\\ \varphi>0\text{ in }\overline{B_{1/2+\rho/2}}\times[-1/2-\rho/2,0],\\ \partial_{t}\varphi-\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}\varphi)<0\text{ in}B_{1+\rho/2}\times(-1-\rho/2,0]\setminus B_{\rho/2}\times(-1-\rho/2,-1+\rho/2].\end{cases}$
Notice that these characteristics resemble the backwards fundamental solution centered at $(0,-1)$ (Figure 25). This problem can be reduced to an ODE assuming that $\varphi$ is a truncation of a function of the form $(t+1)^{-\alpha}\psi(|x|^{2}/(t+1))$.
See [IS12, Lemma 4.16] for further details.
Figure 25. The function $\varphi$ is negative in (a parabolic neighborhood of) $\partial_{\text{par}}(B_{1}\times(-1,0])$ (red region), positive in (a parabolic neighborhood of) $B_{1/2}\times(-1/2,0]$ (yellow region), and satisfies $\partial_{t}\varphi-\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}\varphi)<0$ in (a parabolic neighborhood of) $B_{1}\times(-1,0]\setminus B_{\rho}\times(-1,-1+\rho]$ (complement of the green region). We leave some room so that for every $(y_{0},s_{0})\in B_{\rho/2}\times[0,\rho/2)$, the translations $\varphi_{y_{0},s_{0}}(x,t):=\varphi(x-y_{0},t-s_{0})$ keep similar properties.
4. (Localized Measure Estimate) Follows by combining the previous parts.
1. We can prove it in two steps as in Lemma 3.8 followed by Lemma 3.10.
Let $x_{0}\in A^{1}_{\{0\}}\cap\overline{B_{1/2}}$, and consider for $y_{0}\in B_{1/4}(x_{0})$
$\varphi_{0}^{1}+(\varphi_{y_{0}}^{1}+1/32)=\varphi_{y_{0}/2}^{2}+c_{y_{0}}.$
For $y_{0}\in B_{1/4}(x_{0})$ we have that $\varphi_{y_{0}}^{1}(x_{0})+1/32>0$ and $\varphi_{y_{0}}^{1}+1/32<0$ in ${\mathbb{R}}^{n}\setminus B_{1/2}(x_{0})$ (Figure 26). Therefore
$\emptyset\neq(u-\varphi_{y_{0}/2}^{2})\subseteq B_{1}.$
From this one can show that $|A^{2}_{{\mathbb{R}}^{n}}\cap B_{1}|\geq\eta$.
The localization lemma can be shown modifying the test function from Lemma 3.10. instead of truncating it with a constant we can cap it with a paraboloid that we fit in a $C^{1,1}$ fashion.
See [CC95, Lemma 7.6] for an alternative proof using the ABP maximum principle.
1. (Staking Lemma) We can reduce the problem to the case of zero spatial dimension by using Fubini’s Theorem. Consider then $\mathcal{I}$ to be a collection of disjoint open intervals in $(-1,0)$. For every $I=(t-r,t)$ and $m\geq 1$, let $I^{m}:=(t,t+mr)$. Then we have to show that
$\left|\bigcup_{I\in\mathcal{I}}I^{m}\right|\geq\frac{m}{m+1}\left|\bigcup_{I\in\mathcal{I}}(I^{m}\cup I)\right|.$
Let $T$ be a mapping between intervals such that for $I=(a,b)$, $T(I)=(a-(b-a)/m,b)$. In particular, $T(I^{m})=I^{m}\cup I$. Extend $T$ to unions of disjoint open intervals $E=\bigcup I$ such that $T(E)=\bigcup T(I)$. We still have that the length of $T(E)$ is less than or equal to $(m+1)/m$ times the length of $E$.
On the other hand, any open set of ${\mathbb{R}}$ can we decomposed in a unique way as a numerable union of disjoint open intervals. If $E=\bigcup I$ is a union of non-necessarily disjoint intervals, we get by an inductive argument that $T(E)\supseteq\bigcup T(I)$.
We finally recover the desired estimate by applying the previous observations to the stack $\bigcup_{I\in\mathcal{I}}I^{m}$
$\frac{m+1}{m}\left|\bigcup_{I\in\mathcal{I}}I^{m}\right|\geq\left|T\left(\bigcup_{I\in\mathcal{I}}I^{m}\right)\right|\geq\left|\bigcup_{I\in\mathcal{I}}T(I^{m})\right|=\left|\bigcup_{I\in\mathcal{I}}I^{m}\cup I\right|.$
Figure 26. Assume without loss of generality that $\varphi_{0}^{1}$ touches $u$ from below at $x_{0}\in B_{1/2}$. For $y_{0}\in B_{1/4}(x_{0})$, the paraboloids $\varphi_{y_{0}}^{1}+1/32$ are given such that $u$ must cross the sum $\psi_{y_{0}}:=\varphi_{0}^{1}+(\varphi_{y_{0}}^{1}+1/32)$. Indeed $\psi_{y_{0}}(x_{0})>\varphi_{0}^{1}(x_{0})=u(x_{0})$ and $\psi_{y_{0}}<\varphi_{0}^{1}\leq u$ in ${\mathbb{R}}^{n}\setminus B_{1/2}(x_{0})$.
1. The set $\{m((\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u))_{+}^{n})\geq\delta^{n}\}$ is considered because we need to have that the $L^{n}$-norm of the right-hand side remains small after a rescaling of the form $u_{r}(x)=r^{-2}u(rx+x_{0})$. After this observation, the idea is to use the result from Problem 13.
2. Assume without loss of generality that $\|u\|_{L^{\infty}(B_{3\sqrt{n}})}\leq 1$ and $\|(\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u))_{+}\|_{L^{n}(B_{3\sqrt{n}})}\leq\delta$.
The previous part can be used to show the following diminish of the distribution
$|Q_{1}\setminus A^{M^{k+1}}_{{\mathbb{R}}^{n}}|$ $\leq(1-\eta)|(Q_{1}\setminus A^{M^{k+1}}_{{\mathbb{R}}^{n}})\cup\{m((\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u))_{+}^{n})\geq(\delta M^{k})^{n}\}|$
$\leq(1-\eta)^{k+1}+\sum_{i=0}^{k}(1-\eta)^{k-i}|\{m((\mathcal{P}^{-}_{\lambda,\Lambda}(D^{2}u))_{+}^{n})\geq(\delta M^{i})^{n}\}|$
Finally use the Hardy–Littlewood maximal inequality to bound the distribution of the maximal function in terms of $\delta$.
Details can be consulted in [CC95, Lemma 7.7 and Lemma 7.8].
2. (Weak Harnack) Let $\beta\in C^{\infty}_{c}(B_{5/8})$ radial bump function taking values between 00 and $1$, and such that $\beta=1$ in $B_{1/2}$. Let $\theta\in(0,1]$ such that the graph of $\theta\beta$ touches the graph of $u$ from below at $x_{0}\in B_{5/8}$, and let
$v=u(x_{0})\beta+\mathbbm{1}_{B_{1}\setminus B_{3/4}}u.$
(Figure 27).
Figure 27. The function $v$ is defined piece-wise as a small bump that touches $u$ from below in $B_{5/8}$, and it is equal to $u$ outside $B_{3/4}$. Then the comparison principle between $u$ and $v$ leads to the desired measure estimate.
Then the desired estimate can be completed from the following computation
[…]
3. (Hölder Estimate) The challenge in the implementation of the diminish of oscillation is the non-local aspect of the equation. Even if we assume without loss of generality that $u$ takes values between 00 and $1$ over $B_{1}$, the weak Harnack estimate we showed requires $u$ to be positive in the whole of ${\mathbb{R}}^{n}$.
[…]
We assume in an inductive way and without loss of generality that $u$ satisfies for some small $\alpha,\delta\in(0,1)$
$\begin{cases}|u|\leq(1+\delta)|x|^{\alpha}\text{ in }{\mathbb{R}}^{n}\setminus B_{1},\\ |u|\leq 1\text{ in }B_{1},\\ \|(-\Delta)^{\sigma}u\|_{L^{\infty}(B_{1})}\leq\delta.\end{cases}$
Then we consider whether $u$ is positive or negative at least half of the measure in $B_{1}\setminus B_{3/4}$. If we have that |$|\{u\geq 0\}\cap(B_{1}\setminus B_{3/4})|\geq|B_{1}\setminus B_{3/4}|/2$, then we translate and truncate the function as $v=(u+1)_{+}$. For $\alpha$ sufficiently small we can assure that the right-hand side is also small. Then, by the weak Harnack estimate there is $\theta\in(0,1)$ such that $u+1=v\geq\theta>0$ in $B_{1/2}$.
We need to focus on the relation between the constants $\alpha$, $\delta$, and $\theta$, such as the inductive step holds. Assume that we improved the lower bound as in the previous paragraph. Then we consider $v(x)=(u(x/2)-\theta/2)/(1-\theta/2)$ such that $|v|\leq 1$ in $B_{1}$.
To keep $|v|\leq(1+\delta)|x|^{\alpha}$ in $B_{1}\setminus B_{2}$ we ask that $(2-\theta-2^{1-\alpha})2^{\alpha+1}\geq\theta(2-\theta)$. This can be achieved by choosing $\theta$ sufficiently small in terms of $\alpha$.
Details can be consulted in [Sil06].
1. Follows as in the proof of Corollary 2.10.
1. Follows as in the solution of Problem 6.
1. Follows by combining the weak Harnack inequality and the local maximum principle. Details can be consulted in [CC95, Section 4.2]
1. $v:=-v_{2}$ satisfies $\mathcal{P}^{+}_{\lambda,\Lambda}(D^{2}v)\geq 0$, then by the local maximum principle we just need to provide a bound for $\|v\|_{L^{\varepsilon}(B_{1})}$. This follows by the measure estimate for the Hessian (Problem 15) with a slightly larger exponent.
2. We have to notice that $v_{A}$ is the difference of a convex combination and $u$ and $u$ itself,
[…]
3. Assume by contradiction that for some projection $P_{0}$ it happens that $v_{P_{0}}$ gets smaller than $-(1-\theta)$ for some $\theta\in(0,1)$ to be fixed sufficiently small.
Given that $v_{P_{0}}\geq-1$ satisfies $\mathcal{P}_{\lambda,\Lambda}^{-}(D^{2}v_{P_{0}})\leq 0$, we obtain by the weak Harnack inequality that |$|\{v_{P_{0}}>-1+\theta^{1/2}\}\cap B_{1/4}|\leq C\theta^{\varepsilon/2}$.
Over the complementary set $E:=\{v_{P_{0}}\leq-1+\theta^{1/2}\}\cap B_{1/4}$ set we obtain that
$0\leq\underbrace{\max_{\begin{subarray}{c}P\in{\mathbb{R}}^{n\times n}_{\text{sym}}\\ P^{2}=P\end{subarray}}v_{P}}_{=:v_{+}}-v_{I-P_{0}}=v_{P_{0}}-\underbrace{\min_{\begin{subarray}{c}P\in{\mathbb{R}}^{n\times n}_{\text{sym}}\\ P^{2}=P\end{subarray}}v_{P}}_{=:-v_{-}}\leq\theta^{1/2}$
By uniform ellipticity and $F(0)=0$,
$0\leq\mathcal{P}^{+}_{\lambda,\Lambda}(D^{2}u)=\Lambda v_{+}-\lambda v_{-}.$
Then for $\theta$ sufficiently small we get that $v_{I-P_{0}}\geq\lambda/(2\Lambda)$ in $E$.
However, $w:=\lambda/(2\Lambda)-v_{I-P_{0}}$ satisfies $\mathcal{P}^{+}_{\lambda,\Lambda}(D^{2}w)\geq 0$ and then by the local maximum principle we obtain the following contradiction (for $\theta$ chosen sufficiently small)
$\lambda/(2\Lambda)=w(0)\leq C\|w_{+}\|_{L^{1}(B_{1/4})}\leq C\theta^{\varepsilon/2}.$
Details can be consulted in [CS10].
1. Details can be consulted in [IS16, Proposition 3.3].